### v2

# Markov chain entropy games and the geometry of their Nash equilibria

Abstract.
Consider the following two-person mixed strategy game of a probabilist against Nature with respect to the parameters $(f,\mathcal{B},\pi)$, where $f$ is a convex function satisfying certain regularity conditions, $\mathcal{B}$ is either the set $\{L_{i}\}_{i=1}^{n}$ or its convex hull with each $L_{i}$ being a Markov infinitesimal generator on a finite state space $\mathcal{X}$ and $\pi$ is a given positive discrete distribution on $\mathcal{X}$. The probabilist chooses a prior measure $\mu$ within the set of probability measures on $\mathcal{B}$ denoted by $\mathcal{P}(\mathcal{B})$ and picks a $L\in\mathcal{B}$ at random according to $\mu$, whereas Nature follows a pure strategy to select $M\in\mathcal{L}(\pi)$, the set of $\pi$-reversible Markov generators on $\mathcal{X}$. Nature pays an amount $D_{f}(M||L)$, the $f$-divergence from $L$ to $M$, to the probabilist. We prove that a mixed strategy Nash equilibrium always exists, as well as a minimax result of the form
$\inf_{M\in\mathcal{L}(\pi)}\sup_{\mu\in\mathcal{P}(\mathcal{B})}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)=\sup_{\mu\in\mathcal{P}(\mathcal{B})}\inf_{M\in\mathcal{L}(\pi)}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL).$
This also contrasts with the pure strategy version of the game where we show a Nash equilibrium may not exist. To find approximately a mixed strategy Nash equilibrium, we propose and develop a simple projected subgradient algorithm that provably converges with a rate of $\mathcal{O}(1/\sqrt{t})$, where $t$ is the number of iterations. In addition, we elucidate the relationships of Nash equilibrium with other seemingly disparate notions such as weighted information centroid, Chebyshev center and Bayes risk. This article generalizes the two-person game of a statistician against Nature developed in [D. Haussler (1997). A general minimax result for relative entropy. IEEE Trans. Inform. Theory 43: 1276–1280; D. Haussler and M. Opper (1997). Mutual information, metric entropy and cumulative relative entropy risk. Ann. Statist. 25(6): 2451-2492; A.A. Gushchin and D.A. Zhdanov (2006). A minimax result for $f$-divergences. In: From Stochastic Calculus to Mathematical Finance. Springer, Berlin, Heidelberg.], and highlights the powerful interplay and synergy between modern Markov chains theory and geometry, information theory, game theory, optimization and mathematical statistics.
AMS 2020 subject classifications: 49J35, 60J27, 60J28, 62B10, 62C20, 90C47, 91A05, 91A68, 94A17, 94A29
Keywords: Markov chains; $f$-divergence; information geometry; information centroid; saddle point; Nash equilibrium; minimax theorem; Chebyshev center; Bayes risk; subgradient; algorithmic game theory
1. Introduction
This paper revolves around a two-person game where a probabilist competes against Nature, in which the game itself is parameterized by $(f,\mathcal{B},\pi)$. These parameters will be introduced in a precise manner in subsequent sections, but we nonetheless briefly describe them in order to motivate the investigation of this game. Here, $f$ denotes a convex function adhering to specific regularity conditions, the set $\mathcal{B}$ is either the collection $\{L_{i}\}_{i=1}^{n}$ or its convex hull, with each $L_{i}$ being a Markov infinitesimal generator (see Section 2) on a finite state space $\mathcal{X}$, and $\pi$ is a given discrete distribution on $\mathcal{X}$. We shall offer an original and in-depth study of two versions of the game, namely a pure strategy game and a mixed strategy game. In the latter setting, the probabilist’s mixed strategy involves selecting a prior measure $\mu$ from the set of probability measures on $\mathcal{B}$, denoted by $\mathcal{P}(\mathcal{B})$, and subsequently choosing a random $L\in\mathcal{B}$ based on $\mu$. In contrast, Nature employs a pure strategy to opt for $M\in\mathcal{L}(\pi)$, the set of $\pi$-reversible Markov generators on $\mathcal{X}$. Nature suffers a loss of $D_{f}(M||L)$ to the probabilist, the $f$-divergence from $L$ to $M$. In the pure strategy version of the game, both probabilist and Nature are only allowed to choose deterministically from their respective strategy set.
In the literature, similar games have been developed and analyzed in the context of probability measures. These games are known as the games of statistician against Nature in Haussler (1997); Haussler and Opper (1997); Gushchin and Zhdanov (2006), which appear naturally in statistical estimation. This statistician against Nature game serves as one major motivation and inspiration for us to develop the game of probabilist against Nature. Note that in the game of statistician against Nature, the statistician loses or pays to Nature an amount that depends on the information divergence of the estimator formed by the statistician and the ground truth chosen by Nature, while in the game of probabilist against Nature, the probabilist is awarded by Nature. We justify the interchange of roles of these two players in the context of reversibility and non-reversibility of Markov processes as follows:
Symmetry is a fundamental concept that permeates many disciplines and plays an ubiquitous role across mathematics, natural sciences as well as arts and culture, see for instance the book of Weyl (1952) for a historical account on the notion of symmetry. In view of this perspective, we postulate that Nature, as a player in this game, chooses only from the $\pi$-reversible strategy set. The fact that Nature selects reversible processes follows the current belief of physicists: for instance, the laws of quantum mechanics or of Newton are all time-reversible Schrödinger (1931). In the context of Markov chain Monte Carlo, many classical algorithms in this area such as the Metropolis-Hastings algorithm, Barker proposal or the overdamped Langevin diffusion are naturally motivated by physics and indeed are reversible Markov processes with respect to the stationary distribution.
On the other hand, in hope of designing Markov chains with improved convergence to equilibrium, many probabilists have resorted to non-reversible Markovian Monte Carlo algorithms with provably accelerated convergence rate or improved behaviour, see for example the work Hwang et al. (1993, 2005); Rosenthal and Rosenthal (2015); Bierkens (2016); Diaconis et al. (2000); Rey-Bellet and Spiliopoulos (2016); Gustafson (1998); Kamatani and Song (2023) and the references therein. As a result we postulate that the probabilist, as a competitor against Nature in this game, picks from a strategy set $\mathcal{B}$ that contains in general an inventory of non-reversible chains. The gain of the probabilist, $D_{f}(M||L)$, can be broadly interpreted as an information-theoretic edge or speedup of non-reversibility over reversibility.
While the above explanations justify the role of the probabilist and Nature in the entropy game, from a mathematical point of view however, these two roles can be safely interchanged, or one may wish to replace the term "probabilist" by "Player A" and "Nature" by "Player B" throughout the entire manuscript.
Our discussion so far naturally spurs a number of interesting questions pertaining to these games: is there a Nash equilibrium (Karlin and Peres, 2017; Maschler et al., 2020)? Is the equilibrium unique if it exists? Is there an efficient algorithm to find such an equilibrium and what is the algorithmic complexity? what is the so-called value of the game? Is there a sequential version of this game?
While the number of questions may seem to be endless, in this paper we aim at addressing several fundamental questions regarding the existence and uniqueness of equilibrium, investigating situations where the equilibrium may or may not exist and developing an algorithm to approximately find an equilibrium efficiently. We shall also elaborate on connections with Bayes risk and Wald’s decision game as in classical mathematical statistics.
In addition to the above considerations, another major motivation of the investigation of the game of probabilist against Nature stems from an interesting question raised by Laurent Miclo, which concerns the possibility of providing game-theoretic interpretations to two Metropolis-type Markov generators, namely $P_{-\infty}$ and $P_{\infty}$, which are defined respectively to be, for $x\neq y\in\mathcal{X}$,
$P_{-\infty}(x,y)$ $:=\min\{L(x,y),L_{\pi}(x,y)\},$
$P_{\infty}(x,y)$ $:=\max\{L(x,y),L_{\pi}(x,y)\},$
where the diagonal terms are such that the row sums are zero for all rows and $L_{\pi}$ is the $\pi$-dual of $L$ (see Section 2). Note that $P_{-\infty}$ is the classical Metropolis-Hastings generator in continuous-time while $P_{\infty}$ has been investigated in a series of work in Diaconis and Miclo (2009); Choi (2020); Choi and Huang (2020). In view of the minimum or maximum that appears in these two generators, it seems natural to seek game-theoretic explanations of these two objects. Indeed, as shown in our Example 3.4 below, when the $f$-divergence $D_{f}$ is chosen to be the total variation distance, $P_{-\infty}$, $P_{\infty}$ and their convex combinations can be understood as (part of) a mixed strategy Nash equilibrium of a Markov chain entropy game of probabilist against Nature, thus offering a possible answer to the question of Miclo.
We summarize our main contributions as follow:
  1. (1)
    Introduce the two-person Markov chain entropy games of a probabilist against Nature. This naturally generalizes the game of a statistican against Nature (Haussler, 1997; Haussler and Opper, 1997; Gushchin and Zhdanov, 2006), where instead of Markov generators the game in these references involves probability measures.
    Given the interdisciplinary nature of this topic, the paper draws upon tools and notions to yield new developments in several subjects and uncover interesting connections across these areas, namely
    • modern Markov chains theory (reversiblizations Miclo (1997); Fill (1991); Paulin (2015); Choi (2020) and Markov chain Monte Carlo algorithms Billera and Diaconis (2001); Diaconis and Miclo (2009))
    • geometry (Chebyshev center (Eldar et al., 2008; Candan, 2020), saddle point and information centroid (Choi and Wolfer, 2023), and information geometry of Markov chains Wolfer and Watanabe (2021))
    • information theory ($f$-divergences (Csiszár, 1972; Sason and Verdú, 2016), information geometry (Amari, 2016; Nielsen, 2022), channel capacity and source coding (Davisson and Leon-Garcia, 1980))
    • game theory (Nash equilibrium, value of the game and algorithmic game theory (Karlin and Peres, 2017; Maschler et al., 2020))
    • optimization (minimax or robust optimization, subgradient algorithm (Beck, 2017))
    • mathematical statistics (Bayes risk, Wald’s decision theory and game (Haussler, 1997; Haussler and Opper, 1997; Gushchin and Zhdanov, 2006))
  2. (2)
    Introduce the notion of weighted information centroid in the context of Markov generators and elucidate its key role in understanding the Nash equilibria of the entropy games. We introduce a notion of weighted information centroid of a sequence of Markov chains, which naturally generalizes the notion of information centroid of Markov chains introduced by the authors in Choi and Wolfer (2023). Our analysis shows that the (mixed strategy) Nash equilibrium is intimately related to the notion of Chebyshev center, which can be interpreted as a specific weighted information centroid. This important observation allows us to analyze the existence and uniqueness of Nash equilibrium in the game.
  3. (3)
    Propose and analyze a simple projected subgradient algorithm to find an approximate Nash equilibrium with a provable convergence rate of $\mathcal{O}(1/\sqrt{t})$. A central question in game theory, in particular algorithmic game theory, lies in developing efficient algorithms for (approximate) computation of the Nash equilibrium. To this end, we propose a simple and easy-to-implement projected subgradient algorithm which utilizes the information geometry of the underlying Markov generators.
The rest of this paper is organized as follow. In Section 2, we begin our paper by introducing various notions and in particular the notion of weighted information centroid, which plays a central role in our subsequent investigation. We proceed to present the main results of the paper, where we first address fundamental questions concerning Chebyshev center, weighted information centroids and minimax values in Section 3, followed by a pure strategy game-theoretic analysis in Section 3.1. The corresponding setting of mixed strategy game is presented in Section 3.2, and several simple yet illustrative examples are given in Section 3.1.1 and Section 3.2.1. The design and analysis of a novel projected subgradient algorithm to find an approximate mixed strategy Nash equilibrium is stated in Section 3.3. Finally, the proofs of the main results are incorporated in Section 4.
1.1. Notations
In this subsection we introduce some commonly used notations throughout the manuscript. For $a,b\in\mathbb{Z}$, we write $\llbracket a,b\rrbracket:=\{a,a+1,\ldots,b-1,b\}$ and $\llbracket n\rrbracket:=\llbracket 1,n\rrbracket$ for $n\in\mathbb{N}$. We denote by $\mathbf{0}$ to be the all-zeros matrix on the finite state space $\mathcal{X}$. For a given function $g(n)$, we say that it is $\mathcal{O}(h(n))$ if there exist constants $C>0$ and $n_{0}$ such that $g(n)\leqslant Ch(n)$ for all $n\geqslant n_{0}$.
2. Preliminaries
Consider a convex function $f:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ satisfying $f(1)=0$, and a given positive discrete distribution $\pi$ on a finite state space $\mathcal{X}$. Define $\mathcal{L}$ as the set of Markov infinitesimal generators on $\mathcal{X}$. These generators correspond to $\mathcal{X}\times\mathcal{X}$ matrices having non-negative off-diagonal elements and zero row sums. A generator $L$ is said to be $\pi$-stationary if $\pi L=0$. Moreover, $\mathcal{L}(\pi)\subset\mathcal{L}$ is the subset of $\pi$-reversible generators, where we recall that a generator $L$ is said to be $\pi$-reversible if it satisfies the detailed balance condition, that is, for all $x\neq y\in\mathcal{X}$, we have $\pi(x)L(x,y)=\pi(y)L(y,x).$
In view of (Jansen and Kurt, 2014, Proposition 1.2), we define the $\pi$-dual of a generator $L\in\mathcal{L}$ to be $L_{\pi}$. For $x\neq y$, the off-diagonal elements of $L_{\pi}$ are given by
$L_{\pi}(x,y)=\frac{\pi(y)}{\pi(x)}L(y,x),$
and the diagonal ones ensure zero row sums for all rows. In the special case when $L$ has $\pi$ as its unique stationary distribution, we then have $L_{\pi}=L^{*}$, the adjoint of $L$ in $\ell^{2}(\pi)$ or its time-reversal. Here the space $\ell^{2}(\pi)$ is the standard weighted $\ell^{2}$ Hilbert space endowed with the inner product $\langle\cdot,\cdot\rangle_{\pi}$ given by, for any functions $g,h:\mathcal{X}\rightarrow\mathbb{R}$,
$\langle g,h\rangle_{\pi}:=\sum_{x\in\mathcal{X}}g(x)h(x)\pi(x).$
Following the definition as in Diaconis and Miclo (2009), for a given target $\pi$ and Markov infinitesimal generators $M,L\in\mathcal{L}$, the $f$-divergence from $L$ to $M$ with respect to $\pi$ is defined as
(2.1) $D_{f}(M||L)=\sum_{x\in\mathcal{X}}\pi(x)\sum_{y\in\mathcal{X}\setminus\{x\}}L(x,y)f\left(\frac{M(x,y)}{L(x,y)}\right),$
where the convention of $00f(a/0):=$ 0 for $a\geqslant 0$ applies. If $f^{*}$ is the convex conjugate of $f$, defined by $f^{*}(t)=tf(1/t)$ for $t>0$, then
$D_{f}(M||L)=D_{f^{*}}(L||M),$
and $f^{*}(1)=0$. When $f$ is self-conjugate, that is, $f^{*}=f$, then the $f$-divergence in (2.1) is symmetric.
For a general generator $L$ not necessarily with $\pi$ as its stationary distribution, we consider projecting $L$ onto $\mathcal{L}(\pi)$ under suitable $f$-divergence $D_{f}$. Following Wolfer and Watanabe (2021); Choi and Wolfer (2023), the notions of $f$-projection and $f^{*}$-projection are defined as:
(2.2) $M^{f}=M^{f}(L,\pi):=\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L),\quad M^{f^{*}}=M^{f^{*}}(L,\pi)=\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(L||M).$
To understand our main results, we now introduce the notion of weighted information centroids of Markov chains. It generalizes the notion of (unweighted) information centroids of Markov chains in Choi and Wolfer (2023). Let $\mathbf{w}=(w_{1},\ldots,w_{n})\in\mathbb{R}^{n}_{+}$ be a weight vector in the probability simplex $\mathcal{S}_{n}$ of $n$ elements, that is,
(2.3) $\mathcal{S}_{n}:=\bigg{\{}\mathbf{w}=(w_{1},\ldots,w_{n})\in\mathbb{R}^{n}_{+};~{}\sum_{i=1}^{n}w_{i}=1\bigg{\}}.$
Note that we denote the simplex by $\mathcal{S}_{n}$ instead of $\mathcal{P}(\llbracket n\rrbracket)$ throughout the manuscript. Given a sequence of Markov generators $\{L_{i}\}_{i=1}^{n}$, where $L_{i}\in\mathcal{L}$ for $i\in\llbracket n\rrbracket$, we define the notions of $\mathbf{w}$-weighted $f$-projection centroid and $f^{*}$-projection centroid to be respectively
$M^{f}_{n}$ $=M^{f}_{n}(\mathbf{w},\{L_{i}\}_{i=1}^{n},\pi):=\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i}),$
$M^{f^{*}}_{n}$ $=M^{f^{*}}_{n}(\mathbf{w},\{L_{i}\}_{i=1}^{n},\pi)=\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\sum_{i=1}^{n}w_{i}D_{f}(L_{i}||M).$
We consider two important special cases: first in the case with $n=1$, the above notions reduce to $M_{1}^{f}=M^{f}$ and $M_{1}^{f^{*}}=M^{f^{*}}$ respectively as introduced in (2.2). In the second special case, let $n\in\mathbb{N}$ and $\{\mathbf{e}_{i}\}_{i=1}^{n}$ be the standard unit vectors. It is obvious to see that $M^{f}_{n}(\mathbf{e}_{i},\{L_{i}\}_{i=1}^{n},\pi)=M^{f}(L_{i},\pi)$ and $M^{f^{*}}_{n}(\mathbf{e}_{i},\{L_{i}\}_{i=1}^{n},\pi)=M^{f^{*}}(L_{i},\pi)$.
Our first main result in this Section establishes existence and uniqueness of $f$ and $f^{*}$-projection centroids under strict convexity of $f$, and its proof is given in Section 4.1.
Theorem 2.1 (Existence and uniqueness of weighted $f$ and $f^{*}$-projection centroids).
Suppose we are given a sequence of Markov generators $\{L_{i}\}_{i=1}^{n}$, where $L_{i}\in\mathcal{L}$ for $i\in\llbracket n\rrbracket$, a $f$-divergence $D_{f}$ generated by a strictly convex $f$ which is assumed to have a derivative at $1$ given by $f^{\prime}(1)=0$, and a weight vector $\mathbf{w}\in\mathcal{S}_{n}$. We further assume that there exists at least one $L_{i}\neq\mathbf{0}$ and at least one $w_{i}>0$ when $L_{i}\neq\mathbf{0}$. A weighted $f$-projection of $D_{f}$ (resp. $f^{*}$-projection of $D_{f^{*}}$) that minimizes the mapping
$\mathcal{L}(\pi)\ni M\mapsto\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})\quad\left(\textrm{resp.}~{}=\sum_{i=1}^{n}w_{i}D_{f^{*}}(L_{i}||M)\right)$
exists and is unique, that we denote by $M^{f}_{n}$.
Remark 2.1.
By replacing $f$ by $f^{*}$ in the main result of Theorem 2.1, we can analogously define $M^{f^{*}}_{n}$. Precisely, a weighted $f^{*}$-projection of $D_{f}$ (resp. $f$-projection of $D_{f^{*}}$) that minimizes the mapping
$\mathcal{L}(\pi)\ni M\mapsto\sum_{i=1}^{n}w_{i}D_{f}(L_{i}||M)\quad\left(\textrm{resp.}~{}=\sum_{i=1}^{n}w_{i}D_{f^{*}}(M||L_{i})\right)$
exists and is unique, that we denote by $M^{f^{*}}_{n}$.
Remark 2.2 (On the convexity of $f$).
We emphasize that throughout this manuscript, unless otherwise specified, $f$ is assumed to be a convex function rather than a strictly convex function. This yields the following interesting consequences. Suppose that $f$ is a convex function with $f(s)=0$ for some $s>0,s\neq 1$. For a given $L\in\mathcal{L}$, we thus have
$D_{f}(sL||L)=0,$
and hence the two generators $sL$ and $L$ are indistinguishable with respect to $D_{f}$. From a transition semigroup perspective, this is justifiable since $P^{t}:=e^{sLt}$, the transition semigroup generated by $sL$, is merely a time-change of $Q^{t}:=e^{Lt}$, the transition semigroup generated by $L$, for $t\geqslant 0$.
Remark 2.3.
We discuss the importance of the additional assumptions in Theorem 2.1 in this Remark. In the first case, if $L_{i}=\mathbf{0}$ for all $i\in\llbracket n\rrbracket$, then obviously for any $M\in\mathcal{L}(\pi)$ we have $D_{f}(M||L_{i})=0$ and hence
$\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})=0.$
As such, this shows the existence of weighted $f$-projection centroids and they are not unique even if $f$ is strictly convex. In the second case, suppose that there exists at least one $L_{i}\neq\mathbf{0}$ and $w_{i}=0$ whenever $L_{i}\neq\mathbf{0}$. This again leads to, for any $M\in\mathcal{L}(\pi)$,
$\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})=0.$
We thus exclude the above two degenerate cases in Theorem 2.1.
Remark 2.4.
We stress that in Theorem 2.1, the given positive distribution $\pi$ is arbitrary and at the same time the generators $L_{i}$ need not admit stationary distribution $\pi$ or even be irreducible in the first place. The purpose of Theorem 2.1 is to consider the joint projection of these generators $L_{i}$ onto the space $\mathcal{L}(\pi)$.
In the second result of this Section, we explicitly calculate the weighted $f$ and $f^{*}$-projection centroids $M^{f}_{n}$ and $M^{f^{*}}_{n}$ under some common $f$-divergences. The proof is deferred to Section 4.2. For these common choices of $f$, as shown in Choi and Wolfer (2023) the power mean reversibilizations $P_{p}$ appear naturally as the corresponding $f$ or $f^{*}$-projection. For $x\neq y\in\mathcal{X}$ and $p\in\mathbb{R}\setminus\{0\}$, the off-diagonal entries of $P_{p}$ are given by
(2.4) $P_{p}(x,y):=\left(\frac{L(x,y)^{p}+L_{\pi}(x,y)^{p}}{2}\right)^{1/p},$
where diagonal entries ensure zero row sums. The limiting cases are defined to be
$P_{0}(x,y)$ $:=\lim_{p\to 0}P_{p}(x,y)=\sqrt{L(x,y)L_{\pi}(x,y)},$
(2.5) $P_{\infty}(x,y)$ $:=\lim_{p\to\infty}P_{p}(x,y)=\max\{L(x,y),L_{\pi}(x,y)\},$
(2.6) $P_{-\infty}(x,y)$ $:=\lim_{p\to-\infty}P_{p}(x,y)=\min\{L(x,y),L_{\pi}(x,y)\}.$
Note that in the special case when the weights are given by $w_{i}=1/n$ for all $i\in\llbracket n\rrbracket$, we recover the results in Choi and Wolfer (2023), as expected.
Theorem 2.2 (Examples of weighted $f$ and $f^{*}$-projection centroids).
Given a sequence of Markov generators $(L_{i})_{i=1}^{n}$, where $L_{i}\in\mathcal{L}$ for $i\in\llbracket n\rrbracket$ and a weight vector $\mathbf{w}\in\mathcal{S}_{n}$. We further assume that there exists at least one $L_{i}\neq\mathbf{0}$ and at least one $w_{i}>0$ when $L_{i}\neq\mathbf{0}$. Recall the power mean reversiblizations $P_{p}$ in (2.4).
  1. (1)
    (weighted $f$ and $f^{*}$-projection centroids under $\alpha$-divergence) Let $f(t)=\frac{t^{\alpha}-\alpha t-(1-\alpha)}{\alpha(\alpha-1)}$ for $\alpha\in\mathbb{R}\backslash\{0,1\}$. The unique $f$-projection centroid $M^{f}_{n}$ is given by, for $x\neq y\in\mathcal{X}$,
    $M^{f}_{n}(x,y)=\left(\sum_{i=1}^{n}w_{i}\left(M^{f}(L_{i},\pi)(x,y)\right)^{1-\alpha}\right)^{1/(1-\alpha)},$
    while the unique $f^{*}$-projection centroid $M^{f^{*}}_{n}$ is given by, for $x\neq y\in\mathcal{X}$,
    $M^{f^{*}}_{n}(x,y)=\left(\sum_{i=1}^{n}w_{i}\left(M^{f^{*}}(L_{i},\pi)(x,y)\right)^{\alpha}\right)^{1/\alpha},$
    where $M^{f},M^{f^{*}}$ are respectively the $P_{1-\alpha},P_{\alpha}$-reversiblization.
  2. (2)
    (weighted $f$ and $f^{*}$-projection centroids under squared Hellinger distance) Let $f(t)=(\sqrt{t}-1)^{2}$. The unique $f$-projection centroid $M^{f}_{n}$ is given by, for $x\neq y\in\mathcal{X}$,
    (2.7) $M^{f}_{n}(x,y)=\left(\sum_{i=1}^{n}w_{i}\sqrt{M^{f}(L_{i},\pi)(x,y)}\right)^{2},$
    while the unique $f^{*}$-projection centroid $M^{f^{*}}_{n}$ is given by, for $x\neq y\in\mathcal{X}$,
    $M^{f^{*}}_{n}(x,y)=\left(\sum_{i=1}^{n}w_{i}\sqrt{M^{f^{*}}(L_{i},\pi)(x,y)}\right)^{2},$
    where $M^{f^{*}}=M^{f}$ is the $P_{1/2}$-reversiblization.
  3. (3)
    ($f$ and $f^{*}$-projection centroids under Kullback-Leibler divergence) Let $f(t)=t\ln t-t+1$. The unique $f$-projection centroid $M^{f}_{n}$ is given by, for $x\neq y\in\mathcal{X}$,
    $M^{f}_{n}(x,y)=\prod_{i=1}^{n}\left(M^{f}(L_{i},\pi)(x,y)\right)^{w_{i}},$
    where $00^{0}:=$ 0 in the expression above, while the unique $f^{*}$-projection centroid $M^{f^{*}}_{n}$ is given by, for $x\neq y\in\mathcal{X}$,
    $M^{f^{*}}_{n}(x,y)=\sum_{i=1}^{n}w_{i}M^{f^{*}}(L_{i},\pi)(x,y),$
    where $M^{f},M^{f^{*}}$ are respectively the $P_{0},P_{1}$-reversiblization.
Remark 2.5.
Item (2) can be considered as a special case of item (1) by taking $\alpha=1/2$. Another important special case of $\alpha$-divergence is the $\chi^{2}$-divergence where we take $\alpha=2$.
3. Main results
In this Section, we gradually introduce and state the main results of this paper. We first state two well-known and useful max-min inequalities for $D_{f}$ and $D_{f^{*}}$. Let $\mathcal{A},\mathcal{B}\subseteq\mathcal{L}$. We have
(3.1) $\sup_{L\in\mathcal{B}}\inf_{M\in\mathcal{A}}D_{f}(M||L)$ $\leqslant\inf_{M\in\mathcal{A}}\sup_{L\in\mathcal{B}}D_{f}(M||L),$
(3.2) $\sup_{L\in\mathcal{B}}\inf_{M\in\mathcal{A}}D_{f}(L||M)$ $\leqslant\inf_{M\in\mathcal{A}}\sup_{L\in\mathcal{B}}D_{f}(L||M).$
In the following, we are primarily interested in projection onto the space $\mathcal{L}(\pi)$, and hence we shall take $\mathcal{A}=\mathcal{L}(\pi)$. We are ready to define various minimax and maximin values:
Definition 3.1 (Minimax and maximin values, saddle point property and saddle point).
Let $\mathcal{B}\subseteq\mathcal{L}$. Define
$\overline{v}^{f}=\overline{v}^{f}(\mathcal{B},\pi)$ $:=\inf_{M\in\mathcal{L}(\pi)}\sup_{L\in\mathcal{B}}D_{f}(M||L),$
$\underline{v}^{f}=\underline{v}^{f}(\mathcal{B},\pi)$ $:=\sup_{L\in\mathcal{B}}\inf_{M\in\mathcal{L}(\pi)}D_{f}(M||L).$
We say that $\overline{v}^{f}$ is the minimax value with respect to $(f,\mathcal{B},\pi)$, and similarly $\underline{v}^{f}$ is the maximin value with respect to $(f,\mathcal{B},\pi)$. In view of (3.1) and (3.2), we have
$\overline{v}^{f}\geqslant\underline{v}^{f},\quad\overline{v}^{f^{*}}\geqslant\underline{v}^{f^{*}}.$
If the equality holds, that is,
$\overline{v}^{f}=\underline{v}^{f},$
we say that the saddle point property is satisfied with respect to $(f,\mathcal{B},\pi)$. In this case, suppose that the pair $(M^{f},L^{f})\in\mathcal{L}(\pi)\times\mathcal{L}$ attains the optimal value, that is,
$\overline{v}^{f}=D_{f}(M^{f}||L^{f}),$
then we say that $(M^{f},L^{f})$ is a saddle point with respect to $(f,\mathcal{B},\pi)$.
As mentioned in the introduction of the manuscript, we assume that we are given a set of Markov generators $\{L_{i}\}_{i=1}^{n}$, where $L_{i}\in\mathcal{L}$ for $i\in\llbracket n\rrbracket$, and in the sequel we shall investigate either =$\mathcal{B}=\{L_{i}\}_{i=1}^{n}$ or the convex hull of $\{L_{i}\}_{i=1}^{n}$. Consider the minimax problem
(3.3) $\inf_{M\in\mathcal{L}(\pi)}\max_{i\in\llbracket n\rrbracket}D_{f}(M||L_{i}).$
As $L\mapsto D_{f}(M||L)$ is convex for fixed $M$ and pointwise maximum of convex functions preserves convexity (Boyd and Vandenberghe, 2004, Section $3.2.3$), the mapping $M\mapsto\max_{i\in\llbracket n\rrbracket}D_{f}(M||L_{i})$ is thus convex. As such the outer minimization is a convex minimization problem over the convex set $\mathcal{L}(\pi)$. Inspired by the reformulation of an analogous minimax problem in the context of probability measures in (Candan, 2020, equation $(3)$), the minimax problem (3.3) can be equivalently casted as the following constrained convex minimization:
(3.4) $&\underset{M\in\mathcal{L}(\pi),r}{\min}&&r\\ &\text{s.t.}&&D_{f}(M||L_{i})\leqslant r,~{}\textrm{for all}\,i\in\llbracket n\rrbracket.\end{aligned}$
This problem can be interpreted geometrically as the Chebyshev center problem in the context of Markov chains. For a given $M\in\mathcal{L}(\pi)$ and $r\geqslant 0$, the set $\{L\in\mathcal{L};~{}D_{f}(M||L)\leqslant r\}$ can be understood as the ball of Markov generators within radius $r$ of the center $M$. The constraint in (3.4) thus entails that this ball encloses the set $\{L_{i}\}_{i=1}^{n}$. As the minimization is with respect to the radius $r$, a so-called Chebyshev center is a center of the minimum radius ball that contains $\{L_{i}\}_{i=1}^{n}$, and its optimal value is referred as the Chebyshev radius. Precisely we define the Chebyshev center and radius as follows:
Definition 3.2 (Chebyshev center and radius).
Consider the minimax problem (3.3) and its reformulation (3.4). An optimizer
$\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\max_{i\in\llbracket n\rrbracket}D_{f}(M||L_{i})$
is called a Chebyshev center with respect to $(f,\{L_{i}\}_{i=1}^{n},\pi)$, and the minimax value $\overline{v}^{f}$ is called the Chebyshev radius with respect to $(f,\{L_{i}\}_{i=1}^{n},\pi)$.
For the problem (3.4), we denote the Lagrangian function $\mathsf{L}:\mathbb{R}_{+}\times\mathcal{L}(\pi)\times\mathbb{R}_{+}^{n}\to\mathbb{R}$ to be
$\mathsf{L}(r,M,\mathbf{w}):=r+\sum_{i=1}^{n}w_{i}(D_{f}(M||L_{i})-r),$
where $\mathbf{w}$ is the associated Lagrange multiplier. Differentiating $\mathsf{L}$ with respect to $r$ and setting it to zero gives $\mathbf{w}\in\mathcal{S}_{n}$, the probability simplex that we introduce in (2.3). The dual problem of (3.4) is now written as
$\max_{\mathbf{w}\in\mathbb{R}_{+}^{n}}\min_{r\geqslant 0,M\in\mathcal{L}(\pi)}\mathsf{L}(r,M,\mathbf{w})$ $=\max_{\mathbf{w}\in\mathcal{S}_{n}}\min_{M\in\mathcal{L}(\pi)}\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})$
(3.5) $=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}||L_{i}),$
where we recall that $M^{f}_{n}=M^{f}_{n}(\mathbf{w},\{L_{i}\}_{i=1}^{n},\pi)$ is the $\mathbf{w}$-weighted information centroid as in Theorem 2.1 in the second equality. Note that by weak duality we thus have $\overline{v}^{f}\geqslant\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}||L_{i})$. Our main result below shows that in fact the strong duality holds and connects various notions introduced earlier on. The proof is stated in Section 4.3.
Theorem 3.1.
  1. (1)
    (Strong duality holds for (3.4)) The strong duality holds for (3.4) and there exists $\mathbf{w}^{f}=\mathbf{w}^{f}(\{L_{i}\}_{i=1}^{n},\pi)=(w^{f}_{i})_{i=1}^{n}\in\mathcal{S}_{n}$ such that the optimal value for the dual problem (3.5) is attained at $\mathbf{w}^{f}$, that is,
    $\overline{v}^{f}=\sum_{i=1}^{n}w^{f}_{i}D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{i}).$
    The primal problem (3.4) optimal value is attained at $(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi),\overline{v}^{f})$, and the optimal value of the problem (3.3) is attained at $(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi),j)$, where $j\in\llbracket n\rrbracket$ is such that $w^{f}_{j}>0$.
  2. (2)
    (Chebyshev center is a weighted information centroid $M^{f}_{n}$) A given pair $(M,r)\in\mathcal{L}(\pi)\times\mathbb{R}_{+}$ minimizes the primal problem (3.4) and a given $\mathbf{w}\in\mathcal{S}_{n}$ maximizes the dual problem (3.5) if and only if they satisfy the following two conditions:
    1. (a)
      (Complementary slackness) For $i\in\llbracket n\rrbracket$, we have
      $D_{f}(M||L_{i})\begin{cases}=r&\text{ if }w_{i}>0,\\ \leqslant r&\text{ if }w_{i}=0.\end{cases}$
      Furthermore this implies $r=\overline{v}^{f}$.
    2. (b)
      $M=M^{f}_{n}(\mathbf{w},\{L_{i}\}_{i=1}^{n},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\mathsf{L}(r,M,\mathbf{w})$.
  3. (3)
    (Concavity of the Lagrangian dual) The mapping
    $\mathcal{S}_{n}\ni\mathbf{w}=(w_{i})_{i=1}^{n}\mapsto\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}(\mathbf{w},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})$
    is concave.
  4. (4)
    (Uniqueness of Chebyshev center under strict convexity of $f$) Suppose that the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1. If both $\mathbf{w}_{1},\mathbf{w}_{2}\in\mathcal{S}_{n}$ maximize the dual problem (3.5), we have
    $M^{f}_{n}(\mathbf{w}_{1},\{L_{i}\}_{i=1}^{n},\pi)=M^{f}_{n}(\mathbf{w}_{2},\{L_{i}\}_{i=1}^{n},\pi).$
    In other words, the Chebyshev center is unique.
  5. (5)
    (Characterization of $\overline{v}^{f}=\underline{v}^{f}$) The saddle point property with respect to $(f,\{L_{i}\}_{i=1}^{n},\pi)$ (Definition 3.1) holds, that is,
    $\overline{v}^{f}=\underline{v}^{f}$
    if and only if there exists $j\in\llbracket n\rrbracket$ and $\mathbf{w}^{f}=(w^{f}_{i})_{i=1}^{n}$ a maximizer of the dual problem (3.5) with $w^{f}_{j}>0$ such that
    $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{j}).$
    In particular, if the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1, then the above statement is equivalent to
    $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)=M^{f}(L_{j},\pi).$
  6. (6)
    (Saddle point) If the saddle point property with respect to $(f,\{L_{i}\}_{i=1}^{n},\pi)$ holds, then
    $(M^{f}(L_{l},\pi),L_{l})$
    is a saddle point with respect to $(f,\{L_{i}\}_{i=1}^{n},\pi)$, where $l=\operatorname*{arg\,max}_{i\in\llbracket n\rrbracket}D_{f}(M^{f}(L_{i},\pi)||L_{i})$.
Remark 3.1.
The concavity of the Lagrangian dual in item (3) plays an important role in developing a projected subgradient algorithm for solving (3.5) in Section 3.3.
In Theorem 3.1, we have been investigating the case where =$\mathcal{B}=\{L_{i}\}_{i=1}^{n}$ in Definition 3.1, which can in fact be generalized to the convex hull of these generators. Precisely, we shall take
$\mathcal{B}=\mathrm{conv}((L_{i})_{i=1}^{n}):=\bigg{\{}\sum_{i=1}^{n}\alpha_{i}L_{i};~{}(\alpha_{i})_{i=1}^{n}\in\mathcal{S}_{n},L_{i}\in\mathcal{L}\text{ for all }i\in\llbracket n\rrbracket\bigg{\}}$
and consider minimax problem of the form
(3.6) $\inf_{M\in\mathcal{L}(\pi)}\sup_{L\in\mathrm{conv}((L_{i})_{i=1}^{n})}D_{f}(M||L)=\overline{v}^{f}(\mathrm{conv}((L_{i})_{i=1}^{n}),\pi).$
An important family of Markov generators that can be written as a convex hull is the family of continuized doubly stochastic Markov generators that we denote by $\mathcal{D}$. We say that $L\in\mathcal{D}\subseteq\mathcal{L}$ if all the row and column sums of $L$ are 00 and $L=P-I$ where $P$ is a Markov matrix and $I$ is the identity matrix on $\mathcal{X}$, and it can be shown that the discrete uniform distribution on $\mathcal{X}$ is the stationary distribution of such $L$. Let $P_{i}$ be a permutation matrix on $\mathcal{X}$, then by the Birkhoff-von Neumann theorem (Horn and Johnson, 2013, Theorem $8.7.2$) we have
(3.7) $\mathcal{D}=\mathrm{conv}((P_{i}-I)_{i=1}^{|\mathcal{X}|!}).$
Another family of Markov generators that can be casted under this framework is the set of uniformizable (see (Keilson, 1979, Chapter $2$))and $\mu$-reversible generators, which are used in finite truncation of countably infinite Markov chains in queueing theory van Dijk et al. (2018), see Example 3.7 below.
Our next proposition shows that the analysis in this case of convex hull of generators can be reduced to the setup of Theorem 3.1. The proof can be found in Section 4.4.
Proposition 3.1 (Reduction to the finite case).
$\overline{v}^{f}(\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)$ $=\overline{v}^{f}(\{L_{i}\}_{i=1}^{n},\pi),$
$\underline{v}^{f}(\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)$ $=\underline{v}^{f}(\{L_{i}\}_{i=1}^{n},\pi).$
The next result collects and combines both Theorem 3.1 and Proposition 3.1. Its proof is omitted as it is simply restating these two results in a single Corollary.
Corollary 3.1.
  1. (1)
    (Attainment of the minimax problem (3.6)) There exists $\mathbf{w}^{f}=\mathbf{w}^{f}(\{L_{i}\}_{i=1}^{n},\pi)=(w^{f}_{i})_{i=1}^{n}\in\mathcal{S}_{n}$ such that the problem (3.6) optimal value is attained at $(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi),L_{j})$, where $j\in\llbracket n\rrbracket$ is such that $w^{f}_{j}>0$.
  2. (2)
    (Characterization of $\overline{v}^{f}=\underline{v}^{f}$) The saddle point property with respect to $(f,\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)$ (Definition 3.1) holds, that is,
    $\overline{v}^{f}(\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)=\underline{v}^{f}(\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)$
    if and only if there exists $j\in\llbracket n\rrbracket$ and $\mathbf{w}^{f}=(w^{f}_{i})_{i=1}^{n}$ a maximizer of the dual problem (3.5) with $w^{f}_{j}>0$ such that
    $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{j}).$
    In particular, if the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1, then the above statement is equivalent to
    $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)=M^{f}(L_{j},\pi).$
  3. (3)
    (Saddle point) If the saddle point property with respect to $(f,\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)$ holds, then
    $(M^{f}(L_{l},\pi),L_{l})$
    is a saddle point with respect to $(f,\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)$, where $l=\operatorname*{arg\,max}_{i\in\llbracket n\rrbracket}D_{f}(M^{f}(L_{i},\pi),L_{i})$.
In particular, the above results hold for the set of continuized doubly stochastic Markov generators $\mathcal{D}$ (3.7) with $L_{i}=P_{i}$ and $n=|\mathcal{X}|!$.
3.1. A pure strategy game-theoretic interpretation
In this subsection, we provide a game-theoretic interpretation of the max-min inequalities and the saddle point property (Definition 3.1) and introduce various game-theoretic notions in this context. Much of the exposition of game theory in this section are drawn from content in (Boyd and Vandenberghe, 2004, Section $5.2.5$ and $5.4.3$) and (Karlin and Peres, 2017, Chapter $2$). Consider the following (two-person non-cooperative zero-sum pure strategy) game of a probabilist against the Nature, with respect to the parameters $(f,\mathcal{B},\pi)$. If Nature chooses $M\in\mathcal{L}(\pi)$, the strategy set of Nature, while the probabilist chooses $L\in\mathcal{B}$, the strategy set of the probabilist, then Nature pays an amount $D_{f}(M||L)$ to the probabilist. As a result, Nature wants to minimize $D_{f}$ while the probabilist seeks to maximize $D_{f}$. The quantity $D_{f}(M||L)$ is known as the payoff function of the game, and it is said to be a pure strategy game as both players are only allowed to choose deterministically from their respective strategy sets. On the other hand, a game is said to be a mixed strategy game if the players are allowed to choose randomly within their respective strategy sets. We shall defer to Section 3.2 to discuss the setting of mixed strategy game. In view of Section 3, we shall consider $\mathcal{B}$, the strategy set of the probabilist, to be either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$. In the sequel, we shall call this a two-person pure strategy game of the probabilist against Nature.
Suppose that Nature chooses its strategy $M$ first, followed by the probabilist, who has the knowledge of the choice of Nature. The probabilist thus seeks to choose $L\in\mathcal{B}$ to maximize the payoff $D_{f}(M||L)$. The resulting payoff is $\sup_{L\in\mathcal{B}}D_{f}(M||L)$, which depends on $M$, the choice of Nature. Nature assumes that the probabilist will choose this strategy, and hence Nature seeks to minimize the worst-case payoff, that is, Nature chooses a strategy in the set
$\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\sup_{L\in\mathcal{B}}D_{f}(M||L),$
known as a minimax strategy of Nature, which yields
$\overline{v}^{f}=\overline{v}^{f}(\mathcal{B},\pi)=\inf_{M\in\mathcal{L}(\pi)}\sup_{L\in\mathcal{B}}D_{f}(M||L),$
the payoff from Nature to the probabilist. $\overline{v}^{f}$ is also known as the minimax value of the game.
Now, suppose that the order of play is reversed: the probabilist chooses $L\in\mathcal{B}$ first, and Nature, with the knowledge of $L$, picks from the strategy set $\mathcal{L}(\pi)$. Following a similar argument as before, if both players follow the optimal strategy, the probabilist chooses a strategy in the set
$\operatorname*{arg\,sup}_{L\in\mathcal{B}}\inf_{M\in\mathcal{L}(\pi)}D_{f}(M||L),$
known as a maximin strategy of the probabilist, which yields
$\underline{v}^{f}=\underline{v}^{f}(\mathcal{B},\pi)=\sup_{L\in\mathcal{B}}\inf_{M\in\mathcal{L}(\pi)}D_{f}(M||L),$
the payoff from Nature to the probabilist. $\underline{v}^{f}$ is known as the maximin value of the game.
The max-min inequalities in (3.1) and (3.2), in the context of this two-person game, state that it is advantageous for a player to play second, or more precisely, to know the strategy of the opponent. The difference $\overline{v}^{f}-\underline{v}^{f}\geqslant 0$ can be interpreted as the advantage conferred to a player in knowing the opponent’s strategy. If the minimax equality holds, that is, if $\overline{v}^{f}=\underline{v}^{f}$, then there is no advantage in playing second or in knowing the strategy of the opponent. We now formally define the notion of pure strategy Nash equilibrium and value of the game as follows:
Definition 3.3 (pure strategy Nash equilibrium, value of the game and optimal strategies).
Consider the two-person pure strategy game of the probabilist against Nature as described above. Suppose that the minimax equality holds, that is, $\overline{v}^{f}=\underline{v}^{f}$. We say that
$v=v(f,\mathcal{B},\pi):=\overline{v}^{f}=\underline{v}^{f}$
is the value of the game. Let $M^{f}$ and $L^{f}$ be respectively a minimax strategy of Nature and a maximin strategy of the probabilist. The pair $(M^{f},L^{f})$, a saddle point with respect to $(f,\mathcal{B},\pi)$, is also called a pure strategy Nash equilibrium with respect to $(f,\mathcal{B},\pi)$. $M^{f}$ is said to be an optimal strategy of Nature while $L^{f}$ is said to be an optimal strategy of the probabilist.
Analogous to Corollary 3.1 where we characterize the saddle point, we characterize the existence and uniqueness of Nash equilibrium in this game in the following result. The proof is deferred to Section 4.5.
Corollary 3.2.
Consider the two-person pure strategy game of the probabilist against Nature as described above. Let $\mathcal{B}$ be either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$.
  1. (1)
    (Characterization of pure strategy Nash equilibrium) A pure strategy Nash equilibrium with respect to $(f,\mathcal{B},\pi)$ exists if and only if there exists $j\in\llbracket n\rrbracket$ and $\mathbf{w}^{f}=(w^{f}_{i})_{i=1}^{n}$ a maximizer of the dual problem (3.5) with $w^{f}_{j}>0$ such that
    $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{j}).$
    In particular, if the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1, then the above statement is equivalent to
    $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)=M^{f}(L_{j},\pi).$
  2. (2)
    (Uniqueness of pure strategy Nash equilibrium) Suppose that the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1 and a pure strategy Nash equilibrium exists, which is given by
    $(M^{f}(L_{l},\pi),L_{l}),$
    where $l\in\operatorname*{arg\,max}_{i\in\llbracket n\rrbracket}D_{f}(M^{f}(L_{i},\pi),L_{i})$. It is unique if and only if the index
    $l=\operatorname*{arg\,max}_{i\in\llbracket n\rrbracket}D_{f}(M^{f}(L_{i},\pi),L_{i})$
    is unique.
3.1.1. Examples
In this Section, we give a few simple yet illustrative examples to demonstrate the theory that we have developed thus far. We shall see that, depending on the parameters of the game $(f,\{L_{i}\}_{i=1}^{n},\pi)$, the associated pure strategy Nash equilibrium or saddle point may or may not exist.
Example 3.1 (An example with the existence of multiple pure strategy Nash equilibria (or saddle points)).
In the first example, we take $n=2$ and consider two generators with $L_{1}=L$ and $L_{2}=L_{\pi}$, where $L\in\mathcal{L}$ and recall that $L_{\pi}$ is the $\pi$-dual of $L$.
In view of the bisection property Choi and Wolfer (2023), we have $D_{f}(M||L_{1})=D_{f}(M||L_{2})$, and hence for any weight $\mathbf{w}\in\mathcal{S}_{2}$, we see that
$M^{f}_{2}(\mathbf{w},\{L_{i}\}_{i=1}^{2},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{i})$
for $i=1,2$. Thus, a pure strategy Nash equilibrium or saddle point, with respect to $(f,\{L_{i}\}_{i=1}^{2},\pi)$, exists, according to Corollary 3.2.
Now, we further assume that the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1. Using the bisection property again, the two pure strategy Nash equilibria are
$(M^{f}(L,\pi),L),\quad(M^{f}(L,\pi),L_{\pi}).$
Example 3.2 (An example with a unique pure strategy Nash equilibrium (or saddle point)).
In the second example, we again take $n=2$ and consider two generators with $L_{1},L_{2}\in\mathcal{L}$. We further assume that $D_{f}$ is the $\alpha$-divergence, and $L_{1},L_{2}$ are chosen to satisfy the condition that
(3.8) $D_{f}(M^{f}(L_{1},\pi)||L_{1})>D_{f}(M^{f}(L_{1},\pi)||L_{2}).$
Applying the Pythagorean identity of the $\alpha$-divergence Choi and Wolfer (2023) to the right hand side above, we thus have
$D_{f}(M^{f}(L_{1},\pi)||L_{1})>D_{f}(M^{f}(L_{1},\pi)||L_{2})$ $=D_{f}(M^{f}(L_{2},\pi)||L_{2})+D_{f}(M^{f}(L_{1},\pi)||M^{f}(L_{2},\pi))$
(3.9) $\geqslant D_{f}(M^{f}(L_{2},\pi)||L_{2}).$
Now, we claim that in this example, the pair $(M^{f}(L_{1},\pi),D_{f}(M^{f}(L_{1},\pi)||L_{1}))\in\mathcal{L}(\pi)\times\mathbb{R}_{+}$ minimizes the primal problem (3.4) and $\mathbf{w}=(1,0)\in\mathcal{S}_{2}$ maximizes the dual problem (3.5). This readily follows from item (2) in Theorem 3.1, where the condition (3.8) ensures that the complementary slackness holds.
Using item (1) in Corollary 3.2, we see that a Nash equilibrium exists, and in view of (3.2) and item (2) in Corollary 3.2, the unique pure strategy Nash equilibrium or saddle point is given by
$(M^{f}(L_{1},\pi),L_{1}).$
Example 3.3 (An example where a pure strategy Nash equilibrium (or saddle point) fails to exist).
In the final example, we again specialize into $n=2$ and take two Markov generators with $L_{1},L_{2}\in\mathcal{L}$. We further assume that $D_{f}$ is the $\alpha$-divergence, and $L_{1},L_{2}$ are chosen to satisfy the conditions that $M^{f}(L_{1},\pi)\neq M^{f}(L_{2},\pi)$ and
(3.10) $D_{f}(M^{f}(L_{1},\pi)||L_{1})=D_{f}(M^{f}(L_{2},\pi)||L_{2}).$
Suppose the contrary that a pure strategy Nash equilibrium exists. In view of Corollary 3.2 and $f$ is strictly convex under $\alpha$-divergence, the only possible candidates for the maximizer of the dual problem is either $\mathbf{w}^{f}=(1,0)$ or $\mathbf{w}^{f}=(0,1)$.
In the former case, we thus have $D_{f}(M^{f}(L_{1},\pi)||L_{1})=r$ and
$r$ $=D_{f}(M^{f}(L_{1},\pi)||L_{1})$
$\geqslant D_{f}(M^{f}(L_{1},\pi)||L_{2})$
$=D_{f}(M^{f}(L_{2},\pi)||L_{2})+D_{f}(M^{f}(L_{1},\pi)||M^{f}(L_{2},\pi))$
$=D_{f}(M^{f}(L_{1},\pi)||L_{1})+D_{f}(M^{f}(L_{1},\pi)||M^{f}(L_{2},\pi)),$
where we use the Pythagorean identity Choi and Wolfer (2023) in the second equality and (3.10) in the last equality. This implies that $D_{f}(M^{f}(L_{1},\pi)||M^{f}(L_{2},\pi))=0$ and hence $M^{f}(L_{1},\pi)=M^{f}(L_{2},\pi)$, which contradicts the assumption. Analogously we can handle the case of $\mathbf{w}^{f}=(0,1)$.
As a result, we see that a pure strategy Nash equilibrium fails to exist. For further discussions on this example, please refer to Example 3.5.
3.2. A mixed strategy game of the probabilist against Nature
Recall that in Section 3.1, we introduce a two-person pure strategy game of the probabilist against Nature. In this Section, we shall study a variant of this game where the probabilist uses a mixed strategy while Nature maintains a pure strategy.
Precisely, consider the following game that we refer to as a two-person mixed strategy game of the probabilist against Nature, with respect to the parameters $(f,\mathcal{B},\pi)$. We denote by $\mathcal{P}(\mathcal{B})$ the set of probability measures on $\mathcal{B}$, where again we recall that $\mathcal{B}$ is either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$. The probabilist first chooses a prior measure $\mu\in\mathcal{P}(\mathcal{B})$ and picks a $L\in\mathcal{B}$ at random according to $\mu$. On the other hand Nature, knowing $\mu$ but not knowing $L$, follows a pure strategy to choose $M\in\mathcal{L}(\pi)$. Afterwards, Nature pays an amount $D_{f}(M||L)$ to the probabilist. This serves as a natural generalization of the game of a statistican against Nature proposed and developed in Haussler (1997); Haussler and Opper (1997); Gushchin and Zhdanov (2006).
With these notions in mind, we now introduce the minimax and maximin value of this game:
Definition 3.4 (Minimax and maximin values in the two-person mixed strategy game).
Consider the two-person mixed strategy game of the probabilist against Nature. The minimax value and the maximin value of this game are defined respectively to be
(3.11) $\overline{V}^{f}=\overline{V}^{f}(\mathcal{B},\pi)$ $:=\inf_{M\in\mathcal{L}(\pi)}\sup_{\mu\in\mathcal{P}(\mathcal{B})}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL),$
(3.12) $\underline{V}^{f}=\underline{V}^{f}(\mathcal{B},\pi)$ $:=\sup_{\mu\in\mathcal{P}(\mathcal{B})}\inf_{M\in\mathcal{L}(\pi)}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL).$
Note that $\overline{V}^{f}\geqslant\underline{V}^{f}$. We can define analogously $\overline{V}^{f^{*}},\underline{V}^{f^{*}}$ by replacing $f$ by $f^{*}$ above.
Next, we define the notions of Bayes risk and Bayes strategy of Nature:
Definition 3.5 (Bayes risk and Bayes strategy).
Given a $\mu\in\mathcal{P}(\mathcal{B})$. We say that a Markov generator $M_{\mu}\in\mathcal{L}(\pi)$ is a Bayes strategy with respect to $\mu$ if the mapping $M\mapsto\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)$ attains its infimum at $M_{\mu}$, that is,
$\inf_{M\in\mathcal{L}(\pi)}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)=\int_{\mathcal{B}}D_{f}(M_{\mu}||L)\,\mu(dL).$
The minimum value $\int_{\mathcal{B}}D_{f}(M_{\mu}||L)\,\mu(dL)$ is said to be the Bayes risk with respect to $\mu$,
Finally, we give the notions of minimax and maximin strategy of this game:
Definition 3.6 (Minimax and maximin strategies in the two-person mixed strategy game).
Consider the two-person mixed strategy game of the probabilist against Nature. A strategy in the set
$\operatorname*{arg\,inf}_{M\in\mathcal{L}(\pi)}\sup_{\mu\in\mathcal{P}(\mathcal{B})}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)$
is known as a minimax strategy of Nature, while a strategy in the set
$\operatorname*{arg\,sup}_{\mu\in\mathcal{P}(\mathcal{B})}\inf_{M\in\mathcal{L}(\pi)}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)$
is known as a maximin strategy of the probabilist.
Before we state the main results of this section, we give an important proposition that connects various minimax and maximin values we have introduced thus far, namely $\overline{v}^{f},\overline{V}^{f},\underline{V}^{f}$ and the dual problem (3.5). The proof is deferred to Section 4.6.
Proposition 3.2.
For any $\mathcal{B}\subseteq\mathcal{L}$, we have
(3.13) $\overline{v}^{f}(\mathcal{B},\pi)\geqslant\overline{V}^{f}(\mathcal{B},\pi)\geqslant\underline{V}^{f}(\mathcal{B},\pi),$
where we recall that $\overline{v}^{f}$ is introduced in Definition 3.1 while $\overline{V}^{f},\underline{V}^{f}$ are defined in Definition 3.4. In particular, if $\mathcal{B}$ is either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$, this leads to
(3.14) $\overline{v}^{f}(\mathcal{B},\pi)\geqslant\overline{V}^{f}(\mathcal{B},\pi)\geqslant\underline{V}^{f}(\mathcal{B},\pi)\geqslant\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}||L_{i}).$
In the setting of (3.14), by Theorem 3.1 item (1), the strong duality holds and hence we have
$\overline{v}^{f}(\mathcal{B},\pi)=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}||L_{i}).$
This forces a minimax equality, that is,
$\overline{V}^{f}(\mathcal{B},\pi)=\underline{V}^{f}(\mathcal{B},\pi).$
In other words, all inequalities are in fact equalities in (3.14). We collect this result and a few others in Corollary 3.1 into the following main results:
Theorem 3.2.
Consider the two-person mixed strategy game of the probabilist against Nature, with respect to the parameters $(f,\mathcal{B},\pi)$, where $\mathcal{B}$ is either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$. Recall that there exists $\mathbf{w}^{f}=\mathbf{w}^{f}(\{L_{i}\}_{i=1}^{n},\pi)=(w^{f}_{i})_{i=1}^{n}\in\mathcal{S}_{n}$ such that the optimal value for the dual problem (3.5) is attained at $\mathbf{w}^{f}$ by Theorem 3.1.
  1. (1)
    (A mixed strategy Nash equilibrium always exists) The minimax equality holds, that is,
    $\overline{V}^{f}(\mathcal{B},\pi)=\underline{V}^{f}(\mathcal{B},\pi)=\sum_{i=1}^{n}w^{f}_{i}D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})=:V.$
    We say that $V$ is the value of the game. The pair $(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi),\mathbf{w}^{f})\in\mathcal{L}(\pi)\times\mathcal{P}(\mathcal{B})$ is a mixed strategy Nash equilibrium with respect to the parameters $(f,\mathcal{B},\pi)$.
  2. (2)
    (Game-theoretic interpretation of Chebyshev center and Chebyshev radius) Recall the definition of Chebyshev center and Chebyshev radius in Definition 3.2. A minimax strategy of Nature is given by $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)$, a weighted information centroid, which is also a Chebyshev center. In view of the complementary slackness in Theorem 3.1, we have
    $V=D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{l}),$
    where the index $l$ satisfies $w^{f}_{l}>0$. In words, the Chebyshev radius is the value of the game $V$.
    If the parameters $(f,\{L_{i}\}_{i=1}^{n},\pi)$ satisfy the assumptions in Theorem 2.1, then $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)$ is the unique minimax strategy.
  3. (3)
    (Bayes risk with respect to $\mathbf{w}^{f}$ is value of the game) $\mathbf{w}^{f}$ is a maximin strategy of the probabilist. $M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)$ is a Bayes strategy with respect to $\mathbf{w}^{f}$, and the Bayes risk with respect to $\mathbf{w}^{f}$ is the value of the game $V$.
Remark 3.2.
It is interesting to note that a mixed strategy Nash equilibrium always exists in the two-person mixed strategy game (Theorem 3.2) while a pure strategy Nash equilibrium may or may not exist in the two-person pure strategy game (Corollary 3.2). For an example of the latter case we recall Example 3.3 and another Example 3.5 below. This is analogous to the classical two-person zero-sum game in the game theory literature where a pure strategy Nash equilibrium may not exist, see (Karlin and Peres, 2017, Chapter $2$).
Remark 3.3 (Game-theoretic consequences of the mixed strategy game).
As a mixed strategy Nash equilibrium always exists, there is no advantage for a player (the probabilist or Nature) to play second, or more generally, to know the strategy of the opponent, in the mixed strategy version of the game.
Remark 3.4 (Another proof of $\overline{V}^{f}(\mathcal{B},\pi)=\underline{V}^{f}(\mathcal{B},\pi)$ via the Sion’s minimax theorem).
One common strategy in establishing minimax results, in particular in the context of source coding and information theory, relies on the Sion’s minimax theorem, see for instance (van Erven and Harremoës, 2014, Theorem $35$). For given $\mu\in\mathcal{P}(\mathcal{B})$ and $M\in\mathcal{L}(\pi)$, the mapping
$(M,\mu)\mapsto\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)$
is clearly concave in $\mu$ and convex in $M$. As we are minimizing over $M\in\mathcal{L}(\pi)$, the set of $\pi$-reversible generator $\mathcal{L}(\pi)$ is convex yet unbounded, and hence it is not a compact subset of $\mathcal{L}$. On the other hand, as $\mathcal{B}$ is either a finite set of Markov generators $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$, the set $\mathcal{P}(\mathcal{B})$ is thus compact. As a result, the Sion’s minimax theorem is readily applicable in this setting which yields
$\overline{V}^{f}(\mathcal{B},\pi)=\underline{V}^{f}(\mathcal{B},\pi).$
The proof that we presented in (3.14) relies on the Lagrangian duality theory as in Theorem 3.1, which naturally gives fine properties concerning the mixed strategy Nash equilibrium and related important objects such as $\mathbf{w}^{f}$ and connects with earlier sections of the manuscript. These properties do not seem to be immediate consequences or corollaries from the Sion’s minimax theorem.
3.2.1. Examples
In this Section, similar to Section 3.1.1, we detail four simple examples to demonstrate the theory concerning the two-person mixed strategy game of the probabilist against Nature.
Example 3.4 (Answering a question of Laurent Miclo: an example with the existence of multiple mixed strategy Nash equilibria).
In the first example, we continue the setting discussed in Example 3.1. That is, we consider two ($n=2$) generators with $L_{1}=L$ and $L_{2}=L_{\pi}$, where $L\in\mathcal{L}$.
Using the bisection property of $D_{f}$ Choi and Wolfer (2023), we have $D_{f}(M||L_{1})=D_{f}(M||L_{2})$, and for any weight $\mathbf{w}\in\mathcal{S}_{2}$, we see that
$M^{f}_{2}(\mathbf{w},\{L_{i}\}_{i=1}^{2},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{i})$
for $i=1,2$. In other words, any pair of the form $(M^{f}(L_{1},\pi),\mathbf{w})$ is a mixed strategy Nash equilibrium with respect to the parameters of this game. Note that this game also has multiple pure strategy Nash equilibria as shown in Example 3.1.
In the special case of considering $f(t)=|t-1|$, the $f$-divergence is the total variation distance. We also recall $P_{-\infty},P_{\infty}$ are introduced in (2.6) and (2.5) respectively. It is shown in Choi and Huang (2020) that any convex combinations of $P_{-\infty}$ and $P_{\infty}$ minimize
$\{aP_{-\infty}+(1-a)P_{\infty};~{}a\in[0,1]\}\subseteq\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{i}).$
As such, based on our earlier analysis, $(P_{\infty},\mathbf{w})$ and $(P_{-\infty},\mathbf{w})$ are mixed strategy Nash equilibria of the game under the total variation distance. This gives a possible answer to a question of Laurent Miclo on game-theoretic interpretations of $P_{-\infty},P_{\infty}$.
If $f$ is strictly convex with $f^{\prime}(1)=0$, for instance the $\alpha$-divergence, then $M^{f}(L_{1},\pi)$ is unique, and hence it is the unique minimax strategy of Nature in the mixed strategy game, while any $\mathbf{w}\in\mathcal{S}_{2}$ is a maximin strategy of the probabilist. $M^{f}(L_{1},\pi)$ is also the unique Bayes strategy with respect to any $\mathbf{w}$. The value of the game is given by $V=D_{f}(M^{f}(L_{1},\pi)||L_{1})$.
Note that $M^{f}(L_{1},\pi)$ is also the unique minimax strategy of Nature in the pure strategy game.
Example 3.5 (An example where a pure strategy Nash equilibrium fails to exist and a mixed strategy Nash equilibrium (always) exists).
In the second example, we continue our investigation in Example 3.3 and take $n=2$ along with two Markov generators where $L_{1},L_{2}\in\mathcal{L}$. We choose $D_{f}$ to be the $\alpha$-divergence, and $L_{1},L_{2}$ are assumed to satisfy the conditions that $M^{f}(L_{1},\pi)\neq M^{f}(L_{2},\pi)$ and
(3.15) $D_{f}(M^{f}(L_{1},\pi)||L_{1})=D_{f}(M^{f}(L_{2},\pi)||L_{2}).$
Recall that in Example 3.3, we have already shown that a pure strategy Nash equilibrium cannot exist with these choices of parameters, and furthermore a maximizer of the dual problem (3.5), which exists, cannot be $\mathbf{w}^{f}=(1,0)$ and $\mathbf{w}^{f}=(0,1)$.
We thus come to the conclusion that $w^{f}_{i}>0$ for $i=1,2$. The weighted information centroid $M^{f}_{2}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{2},\pi)$ is the unique minimax strategy of Nature, and is also the unique Bayes strategy with respect to $\mathbf{w}^{f}$.
Figure 1 demonstrates a possible visualization of this example.
Figure 1. A visualization of Example 3.5. Note that the mixed strategy Nash equilibrium is $(M^{f}_{2}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{2},\pi),\mathbf{w}^{f})$, while a pure strategy Nash equilibrium does not exist. The complementary slackness in Theorem 3.1 ensures that $D_{f}(M^{f}_{2}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{2},\pi)||L_{1})=D_{f}(M^{f}_{2}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{2},\pi)||L_{2})=\overline{v}^{f}=V$, the value of the mixed strategy game. (3.15) also holds in this figure.
Example 3.6 (An example where a pure strategy Nash equilibrium and mixed strategy Nash equilibrium coincides).
In this example, we take $n=3$ Markov generators $L_{i}\in\mathcal{L}$ for $i\in\llbracket 3\rrbracket$, and they are chosen in a special way such that
$D_{f}(M^{f}(L_{2},\pi)||L_{1})=D_{f}(M^{f}(L_{2},\pi)||L_{3})=D_{f}(M^{f}(L_{2},\pi)||L_{2}).$
At the same time, we require that, for $j\in\{1,3\}$,
$D_{f}(M^{f}(L_{2},\pi)||L_{2})>D_{f}(M^{f}(L_{j},\pi)||L_{j}).$
With these parameter choices of the game, using Theorem 3.1 and 3.2 we see that we can take $\mathbf{w}^{f}=(0,1,0)$, and $(M^{f}(L_{2},\pi),\mathbf{w}^{f})$ is a mixed strategy Nash equilibrium while $(D_{f}(M^{f}(L_{2},\pi)||L_{2}),2)$ is a pure strategy Nash equilibrium. Figure 2 provides a visualization in this setting.
Figure 2. A visualization of Example 3.6.
Example 3.7 (An example of uniformizable and $\mu$-reversible generators).
Recall that we have been considering settings where $\mathcal{B}$ is either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$. This covers for instance $\mathcal{D}$, the set of doubly stochastic Markov generators, as discussed in (3.7). The aim of this example is to show that this also includes the set of uniformizable and $\mu$-reversible generators, where $\mu\neq\pi$.
Let $L\in\mathcal{L}(\mu)$. $L$ is said to be $\lambda$-uniformizable if $\max_{x\in\mathcal{X}}|L(x,x)|\leqslant\lambda$, and we denote by $\mathcal{L}_{\lambda}(\mu)\subset\mathcal{L}(\mu)$ to be the set of such Markov generators. Without loss of generality, in this example we let $m=|\mathcal{X}|$ and label the state space $\mathcal{X}=\{1,2,\ldots,m\}$. For a given $\lambda>0$, we now claim that
$\mathcal{L}_{\lambda}(\mu)\subseteq\mathrm{conv}(\{L_{x,y}\}_{x<y,x,y\in\mathcal{X}}\cup\{\mathbf{0}\}),$
where for $x<y\in\mathcal{X}$, $L_{x,y}(x,y):=\lambda\dfrac{m(m-1)}{2}$ while $L_{x,y}(y,x):=\lambda\dfrac{m(m-1)}{2}\dfrac{\mu(x)}{\mu(y)}$ and is zero otherwise for other off-diagonal entries of $L_{x,y}$. We also recall that $\mathbf{0}$ is the all-zero Markov generator. Let $L\in\mathcal{L}_{\lambda}(\mu)$. Making use of the $\mu$-reversibility, we write that
$L=\sum_{x<y}w_{x,y}L_{x,y}+\left(1-\sum_{x<y}w_{x,y}\right)\mathbf{0},$
where $w_{x,y}:=\dfrac{L(x,y)}{\lambda\dfrac{m(m-1)}{2}}$.
As a consequence, various results that have been stated are readily applicable to the set $\mathcal{L}_{\lambda}(\mu)$.
3.3. A projected subgradient algorithm to find an approximate mixed strategy Nash equilibrium (or Chebyshev center)
The aim of this Section is to develop a simple and easy-to-implement projected subgradient method to find an approximate mixed strategy Nash equilibrium. Throughout this Section, we consider the two-person mixed strategy game of the probabilist against Nature with respect to the parameters $(f,\mathcal{B},\pi)$, where $\mathcal{B}$ is either $\{L_{i}\}_{i=1}^{n}$ or its convex hull $\mathrm{conv}((L_{i})_{i=1}^{n})$. Recall that $\mathbf{w}^{f}$ is a maximin strategy of the probabilist, and the pair $(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi),\mathbf{w}^{f})$ is a mixed strategy Nash equilibrium. To find such an equilibrium algorithmically, one crucial step amounts to solving the corresponding dual problem (3.5). If the weighted information centroid admits a closed-form expression, for instance in the case of $\alpha$-divergence or the examples given in Theorem 2.2, then making use of $\mathbf{w}^{f}$ a mixed strategy Nash equilibrium is found.
Instead of solving the maximization problem in (3.5), we reformulate this problem into the following equivalent minimization problem by multiplying by $-1$, in order for us to consider the subgradient (rather than supergradient) method in the optimization literature:
$\min_{\mathbf{w}\in\mathcal{S}_{n}}h(\mathbf{w})$
where
(3.16) $h(\mathbf{w}):=-\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})$
Note that the dependence of $h$ on $(f,\mathcal{B},\pi)$ is suppressed to avoid notational burden in the definition above. In view of Theorem 3.1 item (3), the function $h$ is convex.
The algorithm that we propose is novel from two perspectives. First, no algorithm has been developed in the context of statistician against Nature game (Haussler, 1997; Haussler and Opper, 1997; Gushchin and Zhdanov, 2006) to compute the Nash equilibrium, while in this paper we propose a simple subgradient algorithm to approximately find the mixed strategy equilibrium. Second, our algorithm harnesses on the centroid structure and the subgradient of $h$, which has not been observed in the literature of Chebyshev center computation of probability measures (Eldar et al., 2008; Candan, 2020).
In the first main result of this Section, we identify a subgradient of $h$, and its proof is deferred to Section 4.7:
Theorem 3.3 (Subgradient of $h$ and an upper bound of its $\ell^{2}$-norm).
A subgradient of $h$ at $\mathbf{v}\in\mathcal{S}_{n}$ is given by $\mathbf{g}=\mathbf{g}(\mathbf{v})=(g_{1},g_{2},\ldots,g_{n})\in\mathbb{R}^{n}$, where for $i\in\llbracket n\rrbracket$, we have
(3.17) $g_{i}=D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{n})-D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{i}).$
That is, $\mathbf{g}$ satisfies, for $\mathbf{w},\mathbf{v}\in\mathcal{S}_{n}$,
$h(\mathbf{w})\geqslant h(\mathbf{v})+\sum_{i=1}^{n}g_{i}(w_{i}-v_{i}).$
Moreover, the $\ell^{2}$-norm of $\mathbf{g}(\mathbf{v})$ is bounded above by
(3.18) $\left\lVert\mathbf{g}(\mathbf{v})\right\rVert_{2}^{2}:=\sum_{i=1}^{n}g_{i}^{2}\leqslant n\left(|\mathcal{X}|\sup_{\mathbf{v}\in\mathcal{S}_{n};~{}i\in\llbracket n\rrbracket;~{}L_{i}(x,y)>0}L_{i}(x,y)f\left(\dfrac{M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)(x,y)}{L_{i}(x,y)}\right)\right)^{2}=:B.$
Remark 3.5.
The choice of $L_{n}$ in the first term of $g_{i}$ in (3.17) is in fact completely arbitrary, and we shall see in the proof that it can be replaced by any $L_{l}$ with $l\in\llbracket n\rrbracket$. We thus obtain at least $n$ subgradients of $h$ at a given point $\mathbf{v}\in\mathcal{S}_{n}$.
Remark 3.6.
The upper bound $B$ plays an important role in determining the convergence rate of the projected subgradient algorithm, see Theorem 3.4 and its proof below.
We proceed to develop and analyze a projected subgradient algorithm to solve the dual problem (3.5). Suppose that the number of iterations to be ran is $t$.
For $i=1,2,\ldots,t$, we first update the weights of the current iteration by its subgradient, that is, we set
$\mathbf{v}^{(i)}=\mathbf{w}^{(i-1)}-\eta\mathbf{g}(\mathbf{w}^{(i-1)}),$
where $\eta>0$ is the constant stepsize of the algorithm while we recall that $\mathbf{g}$ is a subgradient of $h$ as in (3.17). In the second step, we project $\mathbf{v}^{(i)}$ onto the probability simplex $\mathcal{S}_{n}$ by computing
$\mathbf{w}^{(i)}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{S}_{n}}\left\lVert\mathbf{w}-\mathbf{v}^{(i)}\right\rVert_{2}^{2}.$
This can be done by utilizing existing efficient projection algorithms (see e.g. Condat (2016)), and we do not further investigate these projection algorithms in this manuscript. We then repeat the above two steps for $i=1,2,\ldots,t$. The output of the algorithm is the sequence $(\mathbf{w}^{(i)})_{i=1}^{t}$. Precisely, the algorithm is stated in Algorithm 1.
1
2
Input :  Initial weight value $\mathbf{w}^{(0)}\in\mathcal{S}_{n}$, function $f$ (and $D_{f}$), set $\{L_{i}\}_{i=1}^{n}$, target distribution $\pi$, stepsize $\eta>0$ and number of iterations $t$.
Output :  The sequence $(\mathbf{w}^{(i)})_{i=1}^{t}$.
3
4for $i=1,2,\ldots,t$ do
       // Update via subgradient
5       $\mathbf{v}^{(i)}=\mathbf{w}^{(i-1)}-\eta\mathbf{g}(\mathbf{w}^{(i-1)})$;
       // Projection onto $\mathcal{S}_{n}$
6       $\mathbf{w}^{(i)}=\operatorname*{arg\,min}_{\mathbf{w}\in\mathcal{S}_{n}}\left\lVert\mathbf{w}-\mathbf{v}^{(i)}\right\rVert_{2}^{2}.$
Algorithm 1 A projected subgradient algorithm to find an approximate mixed strategy Nash equilibrium
Before we proceed to analyze the convergence of Algorithm 1, let us develop an intuition on why Algorithm 1 possibly works. Suppose that we are in the hypothetical setting where the optimal weights are all positive, that is, $w^{f}_{i}>0$ for all $i\in\llbracket n\rrbracket$. Recall that by the complementary slackness condition in Theorem 3.1, in this setting the subgradient of $h$ at $\mathbf{w}^{f}$ is zero since
$g_{i}(\mathbf{w}^{f})=D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{j}\}_{j=1}^{n},\pi)||L_{n})-D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})=0.$
As a result, once Algorithm 1 reaches $\mathbf{w}^{f}$, it stays put and no longer moves again, which is a desirable behaviour as an optimal value has been reached. On the other hand, if at the current iteration the subgradient is $g_{i}>0$ at some $i$, then the weights are updated along the subgradient direction followed by projection onto $\mathcal{S}_{n}$.
The success of Algorithm 1 relies crucially on computation of the subgradient $\mathbf{g}(\mathbf{w})$ of $h$ at $\mathbf{w}$, which further depends on the expression $M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)$. For example, it is known in the case of $\alpha$-divergence or the examples given in Theorem 2.2, and hence the algorithm can be readily applied in these settings. Note that in the implementation of Algorithm 1 we assume that the centroid $M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)$ is accessible and we do not discuss algorithms to compute these centroids numerically.
In this paper, we do not consider possibly many variants of the projected subgradient algorithm in Algorithm 1, for instance algorithms with changing or adaptive stepsize, or algorithms with stochastic subgradients.
Our second main result of this Section gives a convergence rate of $\mathcal{O}(1/\sqrt{t})$ of Algorithm 1 with an appropriate choice of stepsize. The proof can be found in Section 4.8. We also write
$\overline{\mathbf{w}}^{t}:=\dfrac{1}{t}\sum_{i=1}^{t}\mathbf{w}^{(i)},$
the arithmetic average up to iteration $t$ of the outputs of Algorithm 1.
Theorem 3.4.
Consider Algorithm 1 with its output $(\mathbf{w}^{(i)})_{i=1}^{t}$ and the notations therein. We have
(3.19) $h\left(\overline{\mathbf{w}}^{t}\right)-h\left(\mathbf{w}^{f}\right)\leqslant\dfrac{n}{2\eta t}+\dfrac{\eta B}{2},$
where we recall that $h$ is the function defined in (3.16) and $B$ is introduced in (3.18). In particular, if we take the constant stepsize to be
$\eta=\sqrt{\dfrac{n}{tB}},$
we thus have
$h\left(\overline{\mathbf{w}}^{t}\right)-h\left(\mathbf{w}^{f}\right)\leqslant\dfrac{1}{2}\sqrt{\dfrac{nB}{t}}+\dfrac{1}{2}\sqrt{\dfrac{nB}{t}}=\sqrt{\dfrac{nB}{t}}.$
Given an arbitrary $\varepsilon>0$, if we further choose
$t=\left\lceil\dfrac{nB}{\varepsilon^{2}}\right\rceil,$
then we reach an $\varepsilon$-close value to $h\left(\mathbf{w}^{f}\right)$ in the sense that
$h\left(\overline{\mathbf{w}}^{t}\right)-h\left(\mathbf{w}^{f}\right)\leqslant\varepsilon.$
Before we introduce the final main result of this Section, we shall fix a few notations. We shall consider the set of continuized generators that does not stay at the same state in one step, that is, for $i\in\llbracket n\rrbracket$ we consider
(3.20) $L_{i}:=P_{i}-I,$
where $I$ is the identity matrix on $\mathcal{X}$ and $P_{i}$ is a transition matrix with $P_{i}(x,x)=0$ for all $x\in\mathcal{X}$. For two Markov generators $L,M\in\mathcal{L}$, by choosing $f(t)=|t-1|/2$, the total variation distance between $L,M$ is given by
$D_{\mathrm{TV}}(L||M):=D_{f}(L||M)=\dfrac{1}{2}\sum_{x\in\mathcal{X}}\pi(x)\sum_{y\in\mathcal{X}\backslash\{x\}}|L(x,y)-M(x,y)|=D_{\mathrm{TV}}(M||L).$
For the family of $\{L_{i}\}_{i=1}^{n}$ satisfying (3.20), we quantify the convergence of $M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{j}\}_{j=1}^{n},\pi)$ towards $M^{f}_{n}(\mathbf{w}^{f},\{L_{j}\}_{j=1}^{n},\pi)$ in terms of the total variation distance $D_{\mathrm{TV}}$ and a strictly convex $f$. The proof can be found in Section 4.9.
Theorem 3.5 ($\mathcal{O}\left(1/\sqrt{t}\right)$ convergence rate of the weighted information centroid generated by Algorithm 1).
Consider Algorithm 1 applied to the family $\{L_{j}\}_{j=1}^{n}$ satisfying (3.20) under a strictly convex $f$ with its output $(\mathbf{w}^{(i)})_{i=1}^{t}$ and the notations therein, and the stepsize is chosen to be
$\eta=\sqrt{\dfrac{n}{tB}}.$
We have
(3.21) $D_{\mathrm{TV}}(M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{j}\}_{j=1}^{n},\pi)||M^{f}_{n}(\mathbf{w}^{f},\{L_{j}\}_{j=1}^{n},\pi))=\mathcal{O}\left(\dfrac{1}{\sqrt{t}}\right),$
where the constant in front on the right hand side depends on all the parameters except $t$, that is, the constant depends on $(f,\{L_{j}\}_{j=1}^{n},\pi,n,\mathcal{X})$.
4. Proofs of the main results
4.1. Proof of Theorem 2.1
The proof is a generalization of (Diaconis and Miclo, 2009, Proposition $1.5$) and (Choi and Wolfer, 2023, Theorem $3.9$) to the setting of weighted information centroid. Pick an arbitrary total ordering on $\mathcal{X}$ with strict inequality being denoted by $\prec$. For $i\in\llbracket n\rrbracket$, we write
$a$ $=a(x,y)=\pi(x)M(x,y),\quad a^{\prime}=a^{\prime}(y,x)=\pi(y)M(y,x),$
$\beta_{i}$ $=\beta_{i}(x,y)=\pi(x)L_{i}(x,y),\quad\beta_{i}^{\prime}=\beta_{i}^{\prime}(y,x)=\pi(y)L_{i}(y,x).$
We note that $M\in\mathcal{L}(\pi)$ gives $a=a^{\prime}$. Using this, we see that
$\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})$ $=\sum_{i=1}^{n}\sum_{x\prec y}w_{i}\pi(x)L_{i}(x,y)f\left(\dfrac{M(x,y)}{L_{i}(x,y)}\right)+w_{i}\pi(y)L_{i}(y,x)f\left(\dfrac{M(y,x)}{L_{i}(y,x)}\right)$
$=\sum_{i=1}^{n}\sum_{x\prec y}w_{i}\beta_{i}f\left(\dfrac{a}{\beta_{i}}\right)+w_{i}\beta_{i}^{\prime}f\left(\dfrac{a}{\beta_{i}^{\prime}}\right)$
$=\sum_{x\prec y}\sum_{i=1}^{n}w_{i}\beta_{i}f\left(\dfrac{a}{\beta_{i}}\right)+w_{i}\beta_{i}^{\prime}f\left(\dfrac{a}{\beta_{i}^{\prime}}\right)$
$=\sum_{x\prec y}\sum_{\{i;~{}w_{i}>0~{}\textrm{and}~{}\beta_{i}>0~{}\textrm{or}~{}\beta_{i}^{\prime}>0\}}w_{i}\beta_{i}f\left(\dfrac{a}{\beta_{i}}\right)+w_{i}\beta_{i}^{\prime}f\left(\dfrac{a}{\beta_{i}^{\prime}}\right)$
$=:\sum_{x\prec y}\Phi_{\mathbf{w},\beta_{1},\ldots,\beta_{n},\beta_{1}^{\prime},\ldots,\beta_{n}^{\prime}}(a).$
To minimize with respect to $M$, we are led to minimize the summand above $\phi:=\Phi_{\mathbf{w},\beta_{1},\ldots,\beta_{n},\beta_{1}^{\prime},\ldots,\beta_{n}^{\prime}}:\mathbb{R}_{+}\to\mathbb{R}_{+}$, where $\mathbf{w}\in\mathcal{S}_{n}$ and $(\beta_{1},\ldots,\beta_{n},\beta_{1}^{\prime},\ldots,\beta_{n}^{\prime})\in\mathbb{R}_{+}^{2n}$ are assumed to be fixed. The summation in the expression of $\phi$ is non-empty in view of the assumptions in the Theorem.
As $\phi$ and $f$ are convex, we denote by $\phi_{+}^{\prime}$ and $f_{+}^{\prime}$ to be their right derivative respectively. It suffices to show the existence of $a_{*}>0$ such that for all $a\in\mathbb{R}_{+}$,
(4.1) $\phi_{+}^{\prime}(a)=\begin{cases}<0,\quad\textrm{ if }\quad a<a_{*},\\ >0,\quad\textrm{ if }\quad a>a_{*}.\end{cases}$
Now, we compute that for all $a\in\mathbb{R}_{+}$,
$\phi_{+}^{\prime}(a)=\sum_{\{i;~{}\beta_{i}>0~{}\textrm{and}~{}\beta_{i}^{\prime}>0\}}w_{i}f^{\prime}_{+}\left(\dfrac{a}{\beta_{i}}\right)+w_{i}f^{\prime}_{+}\left(\dfrac{a}{\beta_{i}^{\prime}}\right)+\sum_{\{i;~{}\beta_{i}>0~{}\textrm{and}~{}\beta_{i}^{\prime}=0\}}w_{i}f^{\prime}_{+}\left(\dfrac{a}{\beta_{i}}\right)+\sum_{\{i;~{}\beta_{i}=0~{}\textrm{and}~{}\beta_{i}^{\prime}>0\}}w_{i}f^{\prime}_{+}\left(\dfrac{a}{\beta_{i}^{\prime}}\right).$
As $f^{\prime}(1)=0$ and $f$ is strictly convex, for sufficiently small $a>0$ $\phi_{+}^{\prime}(a)<0$ while for sufficiently large $a>0$ $\phi_{+}^{\prime}(a)>0$ and $\phi_{+}^{\prime}$ is increasing, we conclude that there exists a unique $a_{*}>0$ such that (4.1) is satisfied.
If we replace $f$ by $f^{*}$ in the above proof, and by noting that $f^{*}$ is also a strictly convex function with $f^{*}(1)=f^{*\prime}(1)=0$, the existence and uniqueness of $M^{f^{*}}_{n}$ is shown.
4.2. Proof of Theorem 2.2
We shall only prove (2.7) as the rest follows exactly the same computation procedure with different choices of $f$. Pick an arbitrary total ordering on $\mathcal{X}$ with strict inequality being denoted by $\prec$. For $i=1,\ldots,n$, we also write
$a$ $=a(x,y)=\pi(x)M(x,y),\quad a^{\prime}=a^{\prime}(y,x)=\pi(y)M(y,x),$
$\beta_{i}$ $=\beta_{i}(x,y)=\pi(x)L_{i}(x,y),\quad\beta_{i}^{\prime}=\beta_{i}^{\prime}(y,x)=\pi(y)L_{i}(y,x).$
The $\pi$-reversibility of $M$ yields $a=a^{\prime}$, which leads to
$\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})$ $=\sum_{i=1}^{n}\sum_{x\prec y}w_{i}\left(\pi(x)L_{i}(x,y)f\left(\dfrac{M(x,y)}{L_{i}(x,y)}\right)+\pi(y)L_{i}(y,x)f\left(\dfrac{M(y,x)}{L_{i}(y,x)}\right)\right)$
$=\sum_{i=1}^{n}\sum_{x\prec y}w_{i}\left(a-2\sqrt{a\beta_{i}}+\beta_{i}+a^{\prime}-2\sqrt{a^{\prime}\beta_{i}^{\prime}}+\beta_{i}^{\prime}\right)$
$=\sum_{x\prec y}\sum_{i=1}^{n}w_{i}\left(2a-2\sqrt{a\beta_{i}}-2\sqrt{a\beta_{i}^{\prime}}+\beta_{i}+\beta_{i}^{\prime}\right).$
Next, we aim to minimize each term in the sum, leading us to minimize the strictly convex function of $a$:
$a\mapsto\sum_{i=1}^{n}w_{i}\left(2a-2\sqrt{a\beta_{i}}-2\sqrt{a\beta_{i}^{\prime}}\right).$
Differentiating the above expression with respect to $a$ yields
$M^{f}_{n}(x,y)=\left(\sum_{i=1}^{n}w_{i}\sqrt{M^{f}(L_{i},\pi)(x,y)}\right)^{2}.$
4.3. Proof of Theorem 3.1
[…]
$M$ $=M^{f}(L_{1},\pi),$
$r$ $=\max_{i\in\llbracket n\rrbracket}D_{f}(M||L_{i})+1>D_{f}(M||L_{l}),$
[…]
$w_{i}(D_{f}(M||L_{i})-r)=0,$
which is equivalent to
$D_{f}(M||L_{i})\begin{cases}=r&\text{ if }w_{i}>0,\\ \leqslant r&\text{ if }w_{i}=0.\end{cases}$
We return to show the second part of item (1). By the first part and item (2), we have
$\overline{v}^{f}=\sum_{i=1}^{n}w^{f}_{i}D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})=D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{l}),$
where $l\in\llbracket n\rrbracket$ is such that $w^{f}_{l}>0$.
Next, we prove item (3). As the Lagrangian dual is concave (Boyd and Vandenberghe, 2004, Section $5.1.2$), in our case this amounts to
$\mathbf{w}\mapsto\min_{r\geqslant 0,M\in\mathcal{L}(\pi)}\mathsf{L}(r,M,\mathbf{w})=\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}||L_{i})$
is concave.
We proceed to prove item (4). Using item (1), we see that the pair $(M^{f}_{n}(\mathbf{w}_{1},\{L_{i}\}_{i=1}^{n},\pi),\overline{v}^{f})$ minimizes the primal problem and $\mathbf{w}_{2}=(w_{2,i})_{i=1}^{n}$ maximizes the dual problem, and hence by item (2), this gives
$M^{f}_{n}(\mathbf{w}_{1},\{L_{i}\}_{i=1}^{n},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\mathsf{L}(r,M,\mathbf{w}_{2})=\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}\sum_{i=1}^{n}w_{2,i}D_{f}(M||L_{i}).$
As $f$ is strictly convex, the minimization problem on the right hand side admits a unique minimizer $M^{f}_{n}(\mathbf{w}_{2},\{L_{i}\}_{i=1}^{n},\pi)$ (Theorem 2.1), and hence $M^{f}_{n}(\mathbf{w}_{1},\{L_{i}\}_{i=1}^{n},\pi)=M^{f}_{n}(\mathbf{w}_{2},\{L_{i}\}_{i=1}^{n},\pi)$.
Next, we prove item (5). We first show the sufficiency. Using item (2), we note that
$\overline{v}^{f}=D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{j})=\min_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{j})\leqslant\max_{l\in\llbracket n\rrbracket}\min_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{l})=\underline{v}^{f}.$
This gives $\overline{v}^{f}=\underline{v}^{f}$ since $\overline{v}^{f}\geqslant\underline{v}^{f}$. To show the necessity, let
$j=\operatorname*{arg\,max}_{l\in\llbracket n\rrbracket}\min_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{l}),$
and we take $\mathbf{w}^{f}$ to be the standard unit vector at the $j$-th coordinate, that is, $w^{f}_{j}=1$ and 00 otherwise. Since $\overline{v}^{f}=\underline{v}^{f}$, this choice of $\mathbf{w}^{f}$ maximizes the dual problem (3.5) and
$M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)\in\operatorname*{arg\,min}_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{j})$
clearly holds.
Finally, we prove item (6). As the saddle point property holds we have
$\overline{v}^{f}=\max_{i\in\llbracket n\rrbracket}D_{f}(M^{f}(L_{i},\pi),L_{i}).$
The stated pair in the theorem statement is thus a saddle point.
4.4. Proof of Proposition 3.1
First, as $D_{f}(\cdot||\cdot)$ is jointly convex in its arguments, and maximization of a convex function over a convex hull occur at an extreme point, we see that
$\sup_{L\in\mathrm{conv}((L_{i})_{i=1}^{n})}D_{f}(M||L)=\max_{i\in\llbracket n\rrbracket}D_{f}(M||L_{i}).$
Taking inf on both sides over $M\in\mathcal{L}(\pi)$ gives the desired result for $\overline{v}^{f}$.
For $\underline{v}^{f}$, we note that the mapping
$L\mapsto\inf_{M\in\mathcal{L}(\pi)}D_{f}(M||L)$
is convex as $D_{f}(\cdot||\cdot)$ is jointly convex and partial minimization of convex function preserves convexity (Boyd and Vandenberghe, 2004, Section $3.2.5$). Using again the property that maximization of a convex function over a convex hull occur at an extreme point, we have
$\underline{v}^{f}(\mathrm{conv}((L_{i})_{i=1}^{n}),\pi)=\sup_{L\in\mathrm{conv}((L_{i})_{i=1}^{n})}\inf_{M\in\mathcal{L}(\pi)}D_{f}(M||L)=\max_{i\in\llbracket n\rrbracket}\inf_{M\in\mathcal{L}(\pi)}D_{f}(M||L_{i})=\underline{v}^{f}(\{L_{i}\}_{i=1}^{n},\pi).$
4.5. Proof of Corollary 3.2
For the first item, it is simply a restatement of the saddle point result Corollary 3.1 in the context of the two-person game and Nash equilibrium.
For the second item, as the index $l$ is unique, using the fact that $f$ is strictly convex $M^{f}(L_{l},\pi)$ is unique, and we thus have a unique saddle point and hence Nash equilibrium. For the other direction, suppose that $l_{1},l_{2}\in\operatorname*{arg\,max}_{i\in\llbracket n\rrbracket}D_{f}(M^{f}(L_{i},\pi),L_{i})$, then both $(M^{f}(L_{l_{1}},\pi),L_{l_{1}})$ and $(M^{f}(L_{l_{2}},\pi),L_{l_{2}})$ are saddle points or Nash equilibria.
4.6. Proof of Proposition 3.2
We first prove the inequalities in (3.13). The inequality $\overline{V}^{f}\geqslant\underline{V}^{f}$ is obvious. To see the first inequality, we note that
$\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)\leqslant\sup_{L\in\mathcal{B}}D_{f}(M||L).$
The desired result follows by taking $\sup_{\mu\in\mathcal{P}(\mathcal{B})}$ then $\inf_{M\in\mathcal{L}(\pi)}$.
Next, we prove the inequalities in (3.14), and it only remains to show the rightmost inequality. We observe that
$\underline{V}^{f}(\mathcal{B},\pi)$ $=\sup_{\mu\in\mathcal{P}(\mathcal{B})}\inf_{M\in\mathcal{L}(\pi)}\int_{\mathcal{B}}D_{f}(M||L)\,\mu(dL)$
$\geqslant\max_{\mathbf{w}\in\mathcal{S}_{n}}\inf_{M\in\mathcal{L}(\pi)}\sum_{i=1}^{n}w_{i}D_{f}(M||L_{i})$
$=\max_{\mathbf{w}\in\mathcal{S}_{n}}\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}||L_{i}),$
which completes the proof.
4.7. Proof of Theorem 3.3
First, by the definition of weighted information centroid, we have
$\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})\leqslant\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{i}).$
Multiplying both sides by $-1$ followed by substracting $h(\mathbf{v})$, we obtain
$h(\mathbf{w})-h(\mathbf{v})$ $=-\sum_{i=1}^{n}w_{i}D_{f}(M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})+\sum_{i=1}^{n}v_{i}D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})$
$\geqslant\sum_{i=1}^{n}-D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})(w_{i}-v_{i})$
$=\sum_{i=1}^{n}-D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})(w_{i}-v_{i})+\sum_{i=1}^{n}D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{n})(w_{i}-v_{i})$
$=\sum_{i=1}^{n}g_{i}(w_{i}-v_{i}),$
where the second equality follows from $\mathbf{w},\mathbf{v}\in\mathcal{S}_{n}$.
Next, we proceed to prove (3.18). First we note that for any $i\in\llbracket n\rrbracket$, we have
$D_{f}(M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})$ $=\sum_{x\in\mathcal{X}}\pi(x)\sum_{y\in\mathcal{X}\backslash\{x\}}L_{i}(x,y)f\left(\dfrac{M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)(x,y)}{L_{i}(x,y)}\right)$
$\leqslant|\mathcal{X}|\sup_{\mathbf{v}\in\mathcal{S}_{n};~{}i\in\llbracket n\rrbracket;~{}L_{i}(x,y)>0}L_{i}(x,y)f\left(\dfrac{M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)(x,y)}{L_{i}(x,y)}\right),$
and hence
$\left\lVert\mathbf{g}\right\rVert_{2}^{2}=\sum_{i=1}^{n}g_{i}^{2}$ $\leqslant\sum_{i=1}^{n}\max\bigg{\{}D_{f}(M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)||L_{n})^{2},D_{f}(M^{f}_{n}(\mathbf{w},\{L_{j}\}_{j=1}^{n},\pi)||L_{i})^{2}\bigg{\}}$
$\leqslant n\max_{l\in\llbracket n\rrbracket}D_{f}(M^{f}_{n}(\mathbf{v},\{L_{j}\}_{j=1}^{n},\pi)||L_{l})^{2}\leqslant B.$
4.8. Proof of Theorem 3.4
We first prove (3.19). For $i\in\llbracket t\rrbracket$, we have
$\left\lVert\mathbf{w}^{(i+1)}-\mathbf{w}^{f}\right\rVert_{2}^{2}$ $\leqslant\left\lVert\mathbf{v}^{(i+1)}-\mathbf{w}^{f}\right\rVert_{2}^{2}$
$=\left\lVert\mathbf{w}^{(i)}-\mathbf{w}^{f}-\eta\mathbf{g}(\mathbf{w}^{(i)})\right\rVert_{2}^{2}$
$=\left\lVert\mathbf{w}^{(i)}-\mathbf{w}^{f}\right\rVert_{2}^{2}+\eta^{2}\left\lVert\mathbf{g}(\mathbf{w}^{(i)})\right\rVert_{2}^{2}-2\eta\mathbf{g}(\mathbf{w}^{(i)})\cdot(\mathbf{w}^{(i)}-\mathbf{w}^{f})$
$\leqslant\left\lVert\mathbf{w}^{(i)}-\mathbf{w}^{f}\right\rVert_{2}^{2}+\eta^{2}B-2\eta\mathbf{g}(\mathbf{w}^{(i)})\cdot(\mathbf{w}^{(i)}-\mathbf{w}^{f}),$
where the first inequality follows from the fact that $\mathbf{w}^{(i+1)}$ is the projection of $\mathbf{v}^{(i+1)}$ onto the simplex $\mathcal{S}_{n}$, while the second inequality follows from the definition of $B$ in (3.18). Upon rearranging and using the definition of subgradient, this leads to
$h(\mathbf{w}^{(i)})-h(\mathbf{w}^{f})$ $\leqslant\mathbf{g}(\mathbf{w}^{(i)})\cdot(\mathbf{w}^{(i)}-\mathbf{w}^{f})$
$=\dfrac{1}{2\eta}\left(\left\lVert\mathbf{w}^{(i)}-\mathbf{w}^{f}\right\rVert_{2}^{2}-\left\lVert\mathbf{w}^{(i+1)}-\mathbf{w}^{f}\right\rVert_{2}^{2}\right)+\dfrac{\eta B}{2}.$
Now, we sum $i$ from $1$ to $t$ to give
$\sum_{i=1}^{t}h(\mathbf{w}^{(i)})-h(\mathbf{w}^{f})$ $\leqslant\dfrac{1}{2\eta}\left(\left\lVert\mathbf{w}^{(1)}-\mathbf{w}^{f}\right\rVert_{2}^{2}-\left\lVert\mathbf{w}^{(t+1)}-\mathbf{w}^{f}\right\rVert_{2}^{2}\right)+\dfrac{\eta Bt}{2}$
$\leqslant\dfrac{1}{2\eta}\left\lVert\mathbf{w}^{(1)}-\mathbf{w}^{f}\right\rVert_{2}^{2}+\dfrac{\eta Bt}{2}\leqslant\dfrac{1}{2\eta}n+\dfrac{\eta Bt}{2},$
where in the last inequality we use the fact that $\mathbf{w}^{(1)},\mathbf{w}^{f}\in\mathcal{S}_{n}$. Dividing both sides by $t$ followed by the convexity of $h$ (see (3.16)) yields
$h\left(\dfrac{1}{t}\sum_{i=1}^{t}\mathbf{w}^{(i)}\right)-h\left(\mathbf{w}^{f}\right)\leqslant\dfrac{1}{t}\left(\sum_{i=1}^{t}h(\mathbf{w}^{(i)})-h(\mathbf{w}^{f})\right)\leqslant\dfrac{n}{2\eta t}+\dfrac{\eta B}{2}.$
The right hand side as a function of the stepsize $\eta$ is minimized when we take $\eta=\sqrt{\dfrac{n}{tB}}$.
4.9. Proof of Theorem 3.5
We first state the following lemma in the setting of this Theorem, which is of independent interest. It can be considered as an extension of (Csiszár, 1972, equation $(3.25)$) to finite Markov chains:
Lemma 4.1.
There exists a constant $\infty>C=C(f,\{L_{i}\}_{i=1}^{n},\pi,n,\mathcal{X})>0$ such that
$D_{\mathrm{TV}}$ $(M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{i}\}_{i=1}^{n},\pi)||M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi))$
$\leqslant C\left(\sum_{i=1}^{n}w^{f}_{i}D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})-\sum_{i=1}^{n}\overline{w}^{t}_{i}D_{f}(M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})\right)$
$=C\left(h\left(\overline{\mathbf{w}}^{t}\right)-h\left(\mathbf{w}^{f}\right)\right).$
Once we have Lemma 4.1, the desired result is obtained, since from Theorem 3.4 with the same choice of stepsize in Algorithm 1 we know that
$D_{\mathrm{TV}}(M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{i}\}_{i=1}^{n},\pi)||M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi))\leqslant C\left(h\left(\overline{\mathbf{w}}^{t}\right)-h\left(\mathbf{w}^{f}\right)\right)=\mathcal{O}\left(\dfrac{1}{\sqrt{t}}\right).$
It thus remains to prove Lemma 4.1. In the remaining of this proof, to minimize notational overhead we write
$M^{f}_{n}(\overline{\mathbf{w}}^{t})=M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{i}\}_{i=1}^{n},\pi),\quad M^{f}_{n}(\mathbf{w}^{f})=M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi).$
Using the strict convexity of $f$ and (3.20), according to (Csiszár, 1972, equation $(3.25)$) we have that
$D_{\mathrm{TV}}$ $(M^{f}_{n}(\overline{\mathbf{w}}^{t})||M^{f}_{n}(\mathbf{w}^{f}))$
$\leqslant C\left(\sum_{i=1}^{n}\overline{w}^{t}_{i}D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})-\sum_{i=1}^{n}\overline{w}^{t}_{i}D_{f}(M^{f}_{n}(\overline{\mathbf{w}}^{t},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})\right)$
$\leqslant C\left(\max_{i\in\llbracket n\rrbracket}D_{f}(M^{f}_{n}(\mathbf{w}^{f},\{L_{i}\}_{i=1}^{n},\pi)||L_{i})+h\left(\overline{\mathbf{w}}^{t}\right)\right)$
$=C\left(-h\left(\mathbf{w}^{f}\right)+h\left(\overline{\mathbf{w}}^{t}\right)\right),$
where the last equality follows from the complementary slackness in Theorem 3.1.
Acknowledgements
Michael Choi would like to thank a question raised by Laurent Miclo on possible game-theoretic interpretations of $P_{-\infty}$ and $P_{\infty}$ when he visited Toulouse School of Economics, which leaves him much to ponder on the topics touch upon by this manuscript. For a possible answer to this question, the readers are referred to Example 3.4. Michael Choi acknowledges the financial support of the project “MAPLE: Mechanistic Accelerated Prediction of Protein Secondary Structure via LangEvin Monte Carlo” with grant number 22-5715-P0001 under the NUS Faculty of Science Ministry of Education Tier 1 grant Data for Science and Science for Data collaborative scheme, project NUSREC-HPC-00001 and NUSREC-CLD-00001 for NUS HPC-AI Priority Projects for Research Program, as well as the startup funding of the National University of Singapore with grant number A-0000178-01-00. Part of this research was supported when Geoffrey Wolfer was part of the Special Postdoctoral Researcher Program (SPDR) of RIKEN and by the Japan Society for the Promotion of Science KAKENHI under Grant 23K13024.